### v2

Abstract
The limited availability of labeled data is a major challenge in audiovisual speech emotion recognition (SER). Self-supervised learning approaches have recently been proposed to mitigate the need for labeled data in various applications. This paper proposes the VQ-MAE-AV model, a vector quantized masked autoencoder (MAE) designed for audiovisual speech self-supervised representation learning and applied to SER. Unlike previous approaches, the proposed method employs a self-supervised paradigm based on discrete audio and visual speech representations learned by vector quantized variational autoencoders. A multimodal MAE with self- or cross-attention mechanisms is proposed to fuse the audio and visual speech modalities and to learn local and global representations of the audiovisual speech sequence, which are then used for an SER downstream task. Experimental results show that the proposed approach, which is pre-trained on the VoxCeleb2 database and fine-tuned on standard emotional audiovisual speech datasets, outperforms the state-of-the-art audiovisual SER methods. Extensive ablation experiments are also provided to assess the contribution of the different model components.
keywords:
Self-supervised learning , Masked autoencoder , Audiovisual speech processing, Emotion recognition
[…]
1 Introduction
Emotion recognition [El Ayadi et al., 2011] aims to automatically identify emotions from unimodal or multimodal data. With the increasing prevalence of audiovisual interfaces, there is a growing need for systems that can accurately recognize emotions from audiovisual speech data. The complementarity of the audio and visual modalities has been demonstrated in various tasks [Gan et al., 2020, Gao & Grauman, 2019, Zhao et al., 2019, Sadok et al., 2024], and using them together can lead to more effective emotion recognition systems. According to Mehrabian [2017], only a small proportion of feelings and emotions is conveyed through the words we use in oral communication. The majority of our feelings and emotions is conveyed through tone, voice, facial expressions, and body language. In recent years, advances in artificial intelligence technology and hardware acceleration have enabled a shift towards multimodal processing in the field of emotion recognition [Ramachandram & Taylor, 2017, Tsai et al., 2019, Schoneveld et al., 2021]. Supervised learning using large annotated datasets remains the dominating paradigm. However, collecting a substantial amount of emotionally labeled data can be expensive, time-consuming, and sometimes impractical. Most datasets for emotion recognition rely on actors to perform various emotions at specific intensities, which requires a considerable amount of acquisition time and resources. In addition, in-the-wild datasets suffer from annotation issues, where annotators fail to reach a consensus, resulting in a significant amount of unusable data [Busso et al., 2008]. Unsupervised and self-supervised learning (SSL) approaches are natural ways to address these issues and have been explored in recent studies [Bengio et al., 2006, Wang et al., 2021, Dib et al., 2023],[Chen & Rudnicky, 2023]. SSL offers the advantage of high scalability, as the SSL task can be carried out on large amounts of unlabeled speech data [Liu et al., 2022, Zhang et al., 2022]. These methods involve pre-training models with pretext tasks and then fine-tuning them on a smaller set of labeled data for the emotion recognition task [Gong et al., 2022a, Pepino et al., 2021, Jegorova et al., 2023].
(a) Discrete visual tokens.
(b) Discrete audio tokens.
Figure 1: Discrete audio and visual tokens creation: (i) fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1); (ii) discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2).
In this paper, we introduce the VQ-MAE-AV model, a vector quantized (VQ) masked autoencoder (MAE) for audiovisual (AV) speech representation learning applied to audiovisual SER. The overall architecture of the model is illustrated in Fig. 1 and 2. To the best of our knowledge, this is the first multimodal SSL approach based on MAEs [He et al., 2022] that is proposed for audiovisual SER. Moreover, unlike existing multimodal MAEs proposed in other application contexts [Bachmann et al., 2022, Geng et al., 2022], the VQ-MAE-AV model operates on discrete latent representations of the modalities and introduces global tokens that are learned using contrastive learning to capture the overall sequence-wise information from each modality. The VQ-MAE-AV model is pre-trained on the VoxCeleb2 dataset [Chung et al., 2018] and fine-tuned on two standard emotional audiovisual speech datasets. The experimental results demonstrate that, across the two test datasets, the proposed VQ-MAE-AV model outperforms the state-of-the-art audiovisual SER methods. Extensive ablation experiments are also presented to study the impact of different model designs. The code and qualitative results are available at https://samsad35.github.io/VQ-MAE-AV.
Figure 2: VQ-MAE-AV model structure. See the first paragraph of Section 3 for a complete description of the pipeline.
2 Related work
Self-supervised learning approaches
SSL models can be broadly categorized into discriminative and generative approaches [Zhang et al., 2022]. Discriminative SSL focuses on creating pairs or groups of data samples and formulating loss functions that enable the model to distinguish or group these samples, which can later benefit downstream tasks [Chen et al., 2020b, a]. Pretext tasks can consist of solving jigsaw puzzles [Noroozi & Favaro, 2016] or predicting image rotations Gidaris et al. [2018] for example. Contrastive learning has emerged as the predominant paradigm in discriminative SSL [Alayrac et al., 2020, Akbari et al., 2021]. In contrast, generative SSL involves generating or reconstructing segments of unlabeled and potentially corrupted data using an autoencoder model [He et al., 2022, Bao et al., 2021, Xie et al., 2022]. The latent representation produced by the encoder can subsequently be used for downstream tasks.
Masked autoencoder
The present paper focuses on the MAE, an SSL generative model that uses an asymmetric encoder-decoder architecture with input masking [He et al., 2022]. The MAE approach is inspired by masked language modeling [Devlin et al., 2019] and has been successfully applied to image modeling thanks to the development of Vision Transformers (ViT) [Dosovitskiy et al., 2020]. The MAE has recently been adapted for audio using a 2D time-frequency representation [Gong et al., 2022a, Baade et al., 2022, Huang et al., 2022]. In the MAE paradigm, the input is divided into non-overlapping patches, each represented by a token embedding. A proportion of tokens are masked, typically 75% for image/audio modeling and 15% for text modeling, and only the visible tokens are fed to the encoder. The encoder outputs are used to reconstruct the masked tokens through a lightweight decoder. To successfully reconstruct the masked tokens, the encoder must generate a semantic representation, which can then be effectively transferred to downstream tasks [He et al., 2022]. It was recently shown that combining the task of reconstructing masked tokens with contrastive learning can improve the representation learned by an MAE [Huang et al., 2023, Gong et al., 2022a].
Extensions of MAE for sequential and multimodal representation learning
A recent extension of the MAE was presented in [Feichtenhofer et al., 2022, Tong et al., 2022] for modeling image sequences, called Video-MAE. It uses the same architecture as the vanilla MAE [He et al., 2022] but incorporates a masking process from video ViT (ViViT) [Arnab et al., 2021]. Since videos often contain redundant information, particularly in scenes with no motion, the authors propose a cubic masking approach along the temporal axis, combined with a high masking ratio of 90%. Other works have extended the MAE to handle multimodal data [Bachmann et al., 2022, Geng et al., 2022, Gong et al., 2022b]. MultiMAE [Bachmann et al., 2022] encodes a small random subset of visible tokens from multiple modalities (RGB, depth, and semantic images) and is trained to reconstruct the missing ones. M3AE [Geng et al., 2022] is a unified MAE architecture for two input modalities (image and text). The main difference in architecture between M3AE and MultiMAE lies in the architecture of the decoder. The MulitMAE approach introduces individual decoders for each modality, enhancing their fusion by incorporating a cross-attention layer at the outset of each decoder. On the other hand, the M3AE approach uses a single decoder that takes the concatenated tokens from all modalities as input.
Adaptation and improvement of the MAE
In the literature, MAEs are typically trained using the L1 or L2 losses, which can negatively affect the reconstruction quality of the masked tokens, resulting, for instance, in blurred or noisy images or sounds. It has been shown that improving the quality of MAE reconstructions can be beneficial in terms of downstream task performance [He et al., 2022]. Several approaches have been proposed in that sense, which include adding a perceptual loss [Dong et al., 2021] or using discrete representations obtained from VQ generative adversarial networks (VQ-GANs) [Esser et al., 2021] or variational autoencoders (VQ-VAEs) [Van Den Oord et al., 2017] to train the MAE [Li et al., 2023, Sadok et al., 2023]. These works only considered an unimodal setting, while the present paper proposes a multimodal MAE for audiovisual speech representation learning. Recently, a model called CAV-MAE [Gong et al., 2022b] was proposed combining MAEs and contrastive learning to learn a representation from audiovisual data. However, different from the proposed VQ-MAE-AV model, CAV-MAE processes raw audiovisual data instead of vector-quantized latent representations, it uses simple mean pooling strategies instead of attention pooling, and it is not applied to SER.
3 The VQ-MAE-AV model
This section presents the VQ-MAE-AV model. The overall approach is illustrated in Fig. 1 and 2 and it can be summarized as follows:
  • 1.
    Fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1);
  • 2.
    Discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2);
  • 3.
    A proportion of the discrete audio and visual tokens is masked out, using a coupled masking strategy between the two modalities (see Section 3.3);
  • 4.
    The visible audio and visual tokens are replaced with trainable continuous embedding vectors (see Section 3.4), which are fed to the VQ-MAE-AV encoder (see Sections 3.5.1 and 3.5.2, where we present two strategies based on attention mechanisms to fuse the modalities);
  • 5.
    Attention pooling is used to compute global sequence-wise tokens that are specific to each modality (see Section 3.5.3);
  • 6.
    The token-wise representation obtained from the encoder is combined with mask tokens and fed to the VQ-MAE-AV decoder, which tries to reconstruct the original discrete audio and visual tokens (see Section 3.5.4);
  • 7.
    The VQ-MAE-AV model is trained in a self-supervised manner to minimize (i) the cross-entropy loss between the reconstructed and original tokens and (ii) a contrastive loss between the audio and visual global tokens (see Section 3.6);
  • 8.
    After self-supervised learning, the VQ-MAE-AV encoder and attention pooling layers are fine-tuned for supervised audiovisual SER (see Section 3.7).
This section will present each above-listed aspect of the model in more detail.
3.1 Vector quantized variational autoencoder
The proposed multimodal self-supervised approach uses the discrete latent representation of two pre-trained and frozen VQ-VAEs [Van Den Oord et al., 2017]. Specifically, as illustrated in Fig. 1, we use the VQ-VAE-audio and VQ-VAE-visual encoders to obtain compressed and quantized representations of the input speech power spectrogram ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}\in\mathbb{R}^{T_{a}\times D}$ where $T_{a}$ and $D$ correspond to the time and frequency dimensions, and of the input image sequence ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}\in\mathbb{R}^{T_{v}\times H\times W\times C}$ where $T_{v}$, $H$, $W$ and $C$ correspond to the time, height, width, and channel dimensions. The audio and visual quantized representations are denoted by ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$, respectively. Each entry of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ corresponds to the index of a vector in the VQ-VAE codebooks. Notably, ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ retains the time-frequency structure of the original spectrogram, while ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ retains the spatio-temporal structure of the original sequence images. This is because the VQ-VAE-audio and VQ-VAE-visual models are designed to be fully convolutional on the frequency and spatial axes, respectively, and they process the frames within a sequence independently. Therefore, compression occurs along the frequency axis ($D^{\prime}\ll D$) for ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and along the x and y-axes of the image ($H^{\prime}\ll H$, $W^{\prime}\ll W$) for ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$. As shown in Fig. 2 and discussed in the following subsections, the proposed MAE-based self-supervised learning approach operates on these discrete and compressed representations before audiovisual speech reconstruction using the VQ-VAE decoders.
The training procedure of the VQ-VAEs follows the original approach presented in Van Den Oord et al. [2017]. In particular, the VQ-VAE loss functions involve a reconstruction term between the original and reconstructed data, which corresponds to the mean squared error for the visual modality and to the Itakura-Saito divergence for the audio modality [Févotte et al., 2009]. More details are provided in Section 4.1.2.
3.2 Discrete audio and visual tokens
As shown in Fig. 1, the audio and visual quantized representations ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$ from the output of the VQ-VAE encoders are divided into non-overlapping patches to build discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot t_{a})\times(n_{d}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot t_{v})\times(n_{h}\cdot h)\times(n_{w}\cdot w)}$, where $T_{a}=n_{t_{a}}\cdot t_{a}$, $D^{\prime}=n_{d}\cdot d$, $T_{v}=n_{t_{v}}\cdot t_{v}$, $H^{\prime}=n_{h}\cdot h$, and $W^{\prime}=n_{w}\cdot w$. These representations are reshaped to ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$, which are seen as sequences of $n_{t_{a}}\cdot n_{d}$ and $n_{t_{v}}\cdot n_{h}\cdot n_{w}$ tokens of dimension $t_{a}\cdot d$ and $t_{v}\cdot h\cdot w$, respectively.
The use of discrete audio and visual tokens in VQ-MAE-AV has several motivations. Firstly, by dividing the audiovisual data into spatio-spectro-temporal patches, the method could learn to relate audio tokens to visual tokens. For example, the audio tokens are expected to correlate strongly with the visual tokens corresponding to the mouth area [Arnela et al., 2016, Sadok et al., 2024]. The proposed method can potentially learn a representation that captures shared and distinctive information between the two modalities by exploiting their complementarity. Additionally, the use of discrete tokens can reduce the computational cost of the method as it involves working with a reduced representation of the data, which allows us to increase the number of tokens (i.e., manipulate longer audiovisual speech sequences) without exploding in the number of trainable parameters compared to the multimodal MAE in the literature [Bachmann et al., 2022].
3.3 Masking
We apply masking to the audio and visual sequences of tokens. The MAE aims to reconstruct the masked tokens from the visible tokens to learn a semantic representation of the data. The masking strategy impacts the performance of downstream tasks. In the original MAE, the masked tokens are randomly drawn with a target ratio of the entire set of tokens, typically 75%. In multimodal scenarios, the basic strategy would be to do the same for each modality, but it has been shown that it is more interesting to implement a coupled masking strategy between the modalities [Bachmann et al., 2022]. This involves randomly drawing a masking proportion $p_{a}\in[0,1]$ for the audio modality and $p_{v}\in[0,1]$ for the visual modality such that $p_{a}+p_{v}=1$. To do this, a Dirichlet distribution is used: $(p_{a},p_{v})\sim Dir(\alpha_{a},\alpha_{v})$ where $\alpha_{a},\alpha_{v}>0$ are the concentration parameters of the distribution. When $\alpha_{a}=\alpha_{v}=1$, the distribution is uniform overall points in its support (the 1-simplex), and $\alpha_{a}\gg\alpha_{b}$ results in a sampling behavior where most of the tokens are taken from the audio modality, and vice versa. These parameters can be used to reduce the dominance of one modality over the other. This strategy allows the reconstruction of missing information from one modality by relying on another. Therefore, the model will leverage the less masked modality for the reconstruction of the more masked one. For this paper, we set $\alpha_{a}=\alpha_{v}=1$.
3.4 Continuous embedding vectors
The discrete tokens correspond to the indices obtained through the quantization step of the pretrained VQ-VAE encoder. Before being input to the VQ-MAE-AV encoder, these discrete tokens are replaced with trainable continuous embedding vectors taken from an audio codebook in $\mathbb{R}^{k_{a}\times e_{a}}$ and from a visual codebook in $\mathbb{R}^{k_{v}\times e_{v}}$, where $k_{a/v}$ is the number of codes in the codebook and $e_{a/v}$ is the dimension of each code. This is simply achieved by replacing the indices of a discrete token with the corresponding vectors of dimension $e_{a/v}$ in the codebook. After this embedding process, the sequences of discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ in $\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$ are transformed into sequences of continuous tokens ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times({t_{a}}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$.
3.5 VQ-MAE-AV encoder and decoder
3.5.1 Attention Block
The VQ-MAE-AV encoder and decoder are built with multi-head Attention blocks similar to those used in the Vision Transformer (ViT) [Dosovitskiy et al., 2020]. Each block comprises a multi-head attention layer, normalization layers, and a Multi-layer Perceptron (MLP). These layers are interconnected with residual connections, as depicted in Fig. 2, and they will be used to capture the inter and intra-relationships between audio and visual tokens. This attention block is inspired by the attention layer in the original transformer [Vaswani et al., 2017]. To simplify the reading afterward, we denote the attention block by $\texttt{Attention}(Q,V,K)$, where $Q,V,K$ are the query, value, and key, respectively. The self-attention mechanism uses the same input vector for the query, key, and value vectors. In the case of cross-attention, the query and key are different to enable attention across multiple modalities or inputs.
3.5.2 Encoders
We propose two fusion strategies of the audio and visual speech data, resulting in two architectures for the VQ-MAE-AV encoder. The first fusion strategy is called self-attention fusion (SAF). As represented in Fig. 2, this fusion consists of a concatenation of token sequences for the two modalities, followed by $L$ self-attention blocks. This concatenation operates on the first dimension of the variables, i.e., we obtain a sequence of $(n_{t_{a}}\cdot n_{d})+(n_{t_{v}}\cdot n_{h}\cdot n_{w})$ tokens after concatenation.
The second fusion strategy is called cross-attention fusion (CAF). As shown in Fig. 2, the sequences of audio and visual tokens are used separately as the queries of two separate attention blocks, which share the same keys and values corresponding to the concatenation of the modalities. These two cross-attention blocks are then followed by a stack of $L$ self-attention blocks.
For both fusion strategies, the encoder outputs one sequence of tokens for each modality, denoted by ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
3.5.3 Global tokens
In addition to the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, we learn two sequence-wise global tokens, denoted by ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$, which can be thought of as similar to [CLS] tokens [He et al., 2022]. These global modality-specific tokens are introduced to aggregate the spectro-temporal and spatio-temporal information in the two modalities, which can be useful for downstream tasks involving predictions at the sequence level, such as audiovisual SER. The global tokens are computed with attention pooling as follows:
${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}$ $=\texttt{Attention}\left(Q_{(a)},V_{(a)},K_{(a)}\right);$ (1)
${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}$ $=\texttt{Attention}\left(Q_{(v)},V_{(v)},K_{(v)}\right),$ (2)
where $Q_{(a)}=\mathcal{Z}^{(a)}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and $Q_{(v)}=\mathcal{Z}^{(v)}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$ represent respectively trainable audio and visual tokens, as proposed in Touvron et al. [2021], $V_{(a)}=K_{(a)}={\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$, and $V_{(v)}=K_{(v)}={\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
3.5.4 Decoders
The token-wise representation obtained from the encoder is combined with mask tokens and fed to the VQ-MAE-AV decoder along with additional position embeddings, as denoted by ${\color[rgb]{1,0,0}\tilde{\mathbf{z}}^{(a)}}$ and ${\color[rgb]{0,0,1}\tilde{\mathbf{z}}^{(v)}}$ and illustrated in Fig. 2. The mask tokens actually correspond to one single trainable vector as proposed in the original MAE [He et al., 2022]. Similarly as for the encoder, the audio and visual inputs of the VQ-MAE-AV decoder can be fused using either self-attention fusion or cross-attention fusion.
The number of attention blocks $L^{\prime}$ in the decoder is chosen to be lower compared to that of the encoder ($L^{\prime}<L$).
A linear layer is added at the end of the decoder, which maps to the size of the VQ-VAE codebooks. The output of this linear layer corresponds to the logits of the discrete tokens. After applying a softmax operation, we obtain reconstructions ${\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}$ and ${\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}$ of the indices ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ that were provided by the VQ-VAE-audio and VQ-VAE-visual encoders, respectively.
3.6 VQ-MAE-AV loss functions
3.6.1 Generative loss function
To train the VQ-MAE-AV model, we minimize the cross-entropy loss applied only to the masked discrete tokens:
$\mathcal{L}_{rec}=\texttt{cross-entropy}\Biggl{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)},~{}{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)}\Biggl{)}$
[…]
where $\mathbf{x}\left(\Omega_{\mathcal{M}}^{(\cdot)}\right)$ denotes the set of masked tokens in $\mathbf{x}$. Another benefit of manipulating discrete representations for multimodal inputs is the homogeneity of the losses, which does not require balancing the losses between the two modalities.
3.6.2 Contrastive loss function
Building upon the approaches presented in Alayrac et al. [2020], Akbari et al. [2021], the global tokens are learned using noise contrastive estimation. This approach enhances the alignment of audiovisual speech pairs by grouping together embeddings that belong to the same time sequence and separating them from those that do not correspond to the same sequence. This approach involves minimizing the loss function $\mathcal{L}_{NCE}({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$, which is defined by:
$\mathcal{L}_{NCE}(\mathbf{u},\mathbf{v})=$
$-\log\left(\frac{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}}{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}+\sum\limits_{(\mathbf{u}^{\prime},~{}\mathbf{v}^{\prime})\in\mathcal{N}}\exp\Bigl{(}\mathbf{u}^{\prime\top}\mathbf{v}^{\prime}/\tau\Bigl{)}}\right).$ (4)
To form positive pairs $({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$ for both audio and visual modalities, we select corresponding streams from the same temporal location in the video. Conversely, negative pairs $({\color[rgb]{1,0,0}\mathbf{w}^{{}^{\prime}(a)}},~{}{\color[rgb]{0,0,1}\mathbf{w}^{{}^{\prime}(v)}})$ are formed by selecting non-corresponding streams drawn from a set $\mathcal{N}$ of different temporal locations for each batch. The sensitivity of the NCE loss in distinguishing between positive and negative pairs is regulated by a temperature parameter $\tau\in\mathbb{R}$.
3.6.3 Final loss function
The final loss function corresponds to the sum of the generative loss function in equation (3.6.1) and of the contrastive loss function in equation (3.6.2).
(a) Qualitative unmasking results for the audio modality. The spectrogram highlighted in the red box represents the original spectrogram. The two spectrograms on the top right represent the spectrograms masked at 50% and 80%, respectively. The reconstructions using VQ-MAE-AV can be seen directly below these masked spectrograms.
(b) Qualitative unmasking results for the visual modality. The first sequence shows the original video, followed by the next two sequences representing the masked video with a ratio of 50% and its reconstruction using VQ-MAE-AV. The last two sequences represent the masked video with a ratio of 80% and its reconstruction using VQ-MAE-AV.
Figure 3: Quantitative results of the audio reconstruction (a) and visual reconstruction (b) using the VQ-MAE-AV model.
3.7 Fine-tuning for audiovisual SER
After pre-training VQ-MAE-AV on token unmasking, the model is fine-tuned for audiovisual SER. We propose three different approaches for fine-tuning the model. The first two approaches rely on a pooling operation (of the encoder outputs ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) to extract the global audio and visual tokens, which is either a simple mean pooling or the attention pooling discussed in Section 3.5.3. The extracted tokens are then concatenated and passed through a single linear layer, followed by an $argmax$ operation to classify the emotion. The third proposed approach, referred to as Query2Emo and inspired by [Liu et al., 2021], involves cross-attention between all audio and visual tokens (concatenation of ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) as key and value, and the emotion classes represented by trainable embeddings as the query. Query2Emo has a single attention block for both the encoder and decoder. The extracted embeddings are then concatenated and passed through a single linear layer. We adopt the asymmetric loss [Ridnik et al., 2021] for all these approaches between the predicted emotion and the ground-truth emotion.
4 Experiments
This section presents our study’s experimental setup and results, as well as a discussion of their implications. We evaluate the effectiveness of the proposed approach for emotion recognition on two standard audiovisual speech databases and compare it with state-of-the-art methods. Additionally, we perform an ablation study to analyze the impact of various hyperparameters and architectures on the performance of our method.
We present qualitative audio and visual reconstruction results in Fig. 3(a) and Fig. 3(b), respectively. The reconstructions are illustrated for two distinct masking ratios: 50% and 80%. Additional qualitative results are available online (see the webpage address at the end of Section 1). Furthermore, we conducted a study to measure the reconstruction quality at different masking ratios of the audiovisual speech data. This analysis is detailed in Section 4.2 and demonstrates the effectiveness of leveraging multimodality to enhance reconstruction quality.
4.1 Experimental setup
4.1.1 Datasets and preprocessing
Dataset for self-supervised training
To pre-train VQ-MAE-AV, we use the VoxCeleb2 dataset [Chung et al., 2018], which offers a broad range of audiovisual speech data from open-source media, with each video featuring a single speaker. We restricted our dataset use to a subset of around 1000 hours of audiovisual speech, encompassing 2170 different speakers. The test set includes about 100 hours of audiovisual speech data with 117 different speakers.
Data pre-processing
The VQ-VAE-audio and VQ-VAE-visual models are trained on the VoxCeleb2 dataset. The former is trained on short-time Fourier transform (STFT) power spectrograms (${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$), while the latter is trained on sequences of RGB images (${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$) captured at 25 fps, cropped and resized to a resolution of $96\times 96$, and aligned using Face-Alignment [Bulat & Tzimiropoulos, 2017]. To compute the STFT, a Hann window of 64 ms (1024 samples at 16 kHz) and a $68$% overlap are used, resulting in a sample rate of 50 Hz, which is twice the sample rate of the visual modality. This leads to sequences of $D=513$ Fourier coefficients.
Emotional audiovisual speech databases
We fine-tune and evaluate the proposed approaches on two emotional audiovisual speech databases.
  • 1.
    RAVDESS [Livingstone & Russo, 2018]: This English database consists of 1440 videos recorded by 24 professional actors and labeled with eight different emotions (neutral, calm, happy, sad, angry, fearful, disgust, surprised).
  • 2.
    CREMA-D [Cao et al., 2014]: It is an English dataset of 7442 videos recorded by 91 actors. Actors spoke from a selection of 12 sentences. The sentences used one of six emotions (anger, disgust, fear, happy, neutral, and sad) with four different intensities (low, medium, high, and unspecified).
We selected these databases as they are commonly used for the emotion recognition task, and their raw data is accessible. We performed 6-fold and 10-fold cross-validation for the RAVDESS and CREAM-D datasets to ensure a fair comparison with previous works. The speaker identities are different in the fine-tuning and evaluation phases.
(a) Peak Signal-to-Noise Ratio (PSNR in dB) on the y-axis as a function of masking ratio (%) on the x-axis.
(b) Signal-to-Distortion Ratio (SDR in dB) on the y-axis as a function of masking ratio (%) on the x-axis.
Figure 4: Quantitative results of the visual reconstruction (a) and audio reconstruction (b). The line represents the mean value across the test dataset, while the shaded area depicts the standard deviation.
4.1.2 Model architecture
VQ-VAE architectures
The VQ-VAE-audio (respectively VQ-VAE-visual) architecture is symmetrical concerning the encoder and the decoder, with three 1D (respectively 2D) convolution for the encoder or transposed convolution for the decoder layers and a residual convolution layer. The VQ-VAE models process each frame independently with no time dependency. For each speech power spectrogram frame of size $D=513$, the VQ-VAE-audio encoder compresses it into a discrete latent vector (a column of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$) of size $D^{\prime}=64$. For each image frame of size $(H=96,W=96,C=3)$, the VQ-VAE-visual encoder compresses it into a discrete latent representation of size $(H^{\prime}=24,W^{\prime}=24)$. The VQ-VAE-audio and the VQ-VAE-visual codebooks contain, respectively, $k_{a}=256$ and $k_{v}=512$ codes of dimension $e_{a}=8$ and $e_{v}=4$. Such a low dimension is chosen to increase the use of the different codes in the codebook [Yu et al., 2021].
Table 1: Accuracy (%) and F1 score (%) results of audiovisual SER. VQ-MAE-AV is pre-trained with both the generative (3.6.1) and contrastive (3.6.2) loss functions using the Query2Emo pooling strategy. The best scores are in bold, and the second-best scores are underlined.
RAVDESS CREMA-D
Method Accuracy F1 score Method Accuracy F1 score
AV-LSTM [Ghaleb et al., 2019] 65.80 - AV-LSTM [Ghaleb et al., 2019] 72.90 -
MuLT [Tsai et al., 2019] 76.60 77.30 MATER [Ghaleb et al., 2020] 67.20 -
CFN-SR [Fu et al., 2021] 75.76 - RAVER [Goncalves & Busso, 2022] 77.30 -
MATER [Ghaleb et al., 2020] 76.30 - MulT Base [Tran & Soleymani, 2022] 68.87 -
AVT [Chumachenko et al., 2022] 79.20 78.20 MulT Large [Tran & Soleymani, 2022] 70.22 -
MDVAE [Sadok et al., 2024] 79.30 80.70 AV-Gating [Ghaleb et al., 2019] 74.00 -
VQ-MAE-AV (ours) 84.80 84.50 VQ-MAE-AV (ours) 80.40 80.00
Table 2: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the attention pooling strategy for the fine-tuning. ‘Pre-train’ refers to the training of the VQ-MAE-AV for the unmasking task on the VoxCeleb2 database. ‘Freeze’ refers to the freezing of the VQ-MAE-AV encoder.
Method Pre-train Freeze Accuracy (%)
VQ-MAE-AV 29.6
VQ-MAE-AV 70.5
VQ-MAE-AV 81.5
Table 3: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, fine-tuned on RAVDESS for different encoder depths.
Method Parm. (M) Acc. (%) F1 score (%)
VQ-MAE-AV-6 8.5 75.7 76.0
VQ-MAE-AV-12 13.5 81.5 80.1
VQ-MAE-AV-16 16.8 82.4 82.4
VQ-MAE-AV-20 20.2 81.3 81.3
Table 4: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the Query2Emo pooling strategy for the fine-tuning. ‘Generative’ corresponds to the loss function in Eq. 3.6.1 and ‘Contrastive’ corresponds to the loss function in Eq. 3.6.2.
Method Contrastive Generative Accuracy (%)
VQ-MAE-AV 75.2
VQ-MAE-AV 84.3
VQ-MAE-AV 84.8
Table 5: Performance of VQ-MAE-AV without contrastive learning for different encoder and decoder architectures, using the attention pooling strategy for the fine-tuning. SAF stands for self-attention fusion and CAF stands for cross-attention fusion.
Method Param. (M) Encoder Decoder Acc. (%)
VQ-MAE-AV 13.5 SAF SAF 81.5
VQ-MAE-AV 25.0 CAF SAF 82.8
VQ-MAE-AV 18.4 SAF CAF 81.8
VQ-MAE-AV 30.0 CAF CAF 83.0
Table 6: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, without contrastive learning, and using different pooling strategies for the fine-tuning.
Pooling strategy Accuracy (%) F1 score (%)
Mean Pooling 78.1 78.4
Attention Pooling 81.5 80.1
Query2Emo 84.3 84.8
Table 7: Comparison of performance between unimodal VQ-MAE-A-12 or VQ-MAE-V-12 and multimodal VQ-MAE-AV-12 (using SAF for both encoder and decoder and without contrastive learning).
Method Accuracy (%) F1 score (%)
VQ-MAE-A-12 73.2 72.8
VQ-MAE-V-12 74.1 73.9
VQ-MAE-AV-12 81.5 80.1
VQ-MAE-AV architectures
The VQ-MAE-AV model uses $L=12$ attention blocks in the encoder and $L^{\prime}=4$ in the decoder. Each self-attention layer of a block is divided into four heads. By default, the parameters of the discrete audio and visual tokens ($d$, $h$, and $w$) are set to 4. We set $t_{a}=10$ and $t_{v}=5$ because the sampling rate of ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$ is twice the sampling rate of ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$. In the ablation study (Section 4.4), we will explore all possible combinations of the VQ-MAE-AV encoder and decoder, according to the two fusion strategies (self-attention fusion and cross-attention fusion). We will also evaluate the impact of the pooling strategy and contrastive learning.
4.1.3 Training and fine-tuning details
Self-supervised training details
The VQ-MAE-AV is trained using the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate, with a 100-epoch warm-up period. The parameters of the optimizer, similar to He et al. [2022], are $\beta_{2}=0.9$, $\beta_{2}=0.95$, and weight_decay$=0.05$. The base learning rate follows the linear scaling rule [Goyal et al., 2017] $lr=(\texttt{base\_lr}=1e-3)\times(\texttt{batchsize}=128)/256$. We distributed the pre-training of VQ-MAE-AV on 4 NVIDIA HGX A100. The training lasted for 160 epochs, and each epoch took approximately 15 minutes.
Fine-tuning details
For the fine-tuning process, we also use the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate and with a 40-epoch warm-up period. The parameters of the optimizer are the same as those used for the pre-training. The base learning rate is 1e-4.
4.2 Audiovisual speech reconstruction quality
In this experiment, we evaluate the reconstruction quality of VQ-MAE-AV when applied to masked audiovisual speech data. The model is fed with a sequence of tokens, some of which have been masked, and it is used to predict the masked tokens, i.e. to reconstruct the complete audiovisual speech sequence from partially observed tokens. We will study the reconstruction performance for different masking ratios.
For this experiment, we compare VQ-MAE-AV to VQ-MAE-A (A for audio) and VQ-MAE-V (V for visual), which are the unimodal versions of VQ-MAE-AV. The average quality performance for the speech and visual modalities is evaluated using the VoxCeleb2 test set. The Peak Signal-to-Noise Ratio (PSNR in dB) is used to assess the quality of the resynthesized visual data, and the Signal-to-Distortion Ratio (SDR in dB) is used to assess the quality of the resynthesized audio data.
Fig. 4(a) and Fig.4(b) present the PSNR and SDR curves, respectively, for the reconstruction quality of the visual and audio modalities as a function of the masking ratio. Notably, VQ-MAE-AV outperforms VQ-MAE-V for masking ratios greater than $50$%. At a masking ratio of $90$%, VQ-MAE-AV achieves a significant $3.68$ dB gain in PSNR over VQ-MAE-V. For the audio modality, VQ-MAE-AV outperforms VQ-MAE-A for masking ratios greater than $40$% and records a gain of $2.87$ dB in SDR at $90$% of masking. In summary, this experiment highlights the effectiveness of leveraging multimodality to improve reconstruction quality.
4.3 Audiovisual emotion recognition
Table 1 compares the emotion recognition performance (accuracy and F1 score) of the proposed VQ-MAE-AV model (using the cross-attention fusion strategy for both the encoder and decoder) with the performance of several audiovisual state-of-the-art methods.
The experimental comparison includes LSTM-based methods [Ghaleb et al., 2019] and Transformer-based methods [Tsai et al., 2019, Chumachenko et al., 2022, Goncalves & Busso, 2022]. MATER [Ghaleb et al., 2020] uses distinct Transformer architectures for each modality, merging the embeddings via mean pooling. CFN-SR [Fu et al., 2021] introduces a cross-modality learning approach, employing self-attention and residual structures to model inter- and intra-modality interactions. We also consider the AVT method [Chumachenko et al., 2022], which uses self-attention fusion and modality dropout to address the challenge of one modality being absent. Additionally, RAVER [Goncalves & Busso, 2022] addresses challenges related to modality alignment, temporal information capture, and handling missing features.
The experimental comparison also includes unsupervised audiovisual approaches, such as the multimodal dynamical VAE (MDVAE) [Sadok et al., 2024], which employs a hierarchical latent space to segregate static from dynamic information and modality-common from modality-specific information. Additionally, we compare with MulT [Tran & Soleymani, 2022], a self-supervised approach using the Transformer architecture and pre-trained on the VoxCeleb dataset with a pretext task of reconstructing masked frames (with a masking ratio of 15%).
As can be seen in Table 1, the VQ-MAE-AV model with Query2Emo outperforms the most recent methods overall. VQ-MAE-AV achieves $9.04$%, $5.60$% and $5.50$% better accuracy than the CFN-SR [Fu et al., 2021], AVT [Chumachenko et al., 2022] and MDVAE [Sadok et al., 2024] methods for the RAVDESS dataset, respectively. Regarding the CREMA-D dataset, VQ-MAE-AV achieves $7.50$%, $6.40$%, and $3.10$% better accuracy than the AV-LSTM [Ghaleb et al., 2019], AV-Gating [Ghaleb et al., 2019], and RAVER [Goncalves & Busso, 2022]. Overall, the experimental results demonstrate the effectiveness of the proposed audiovisual self-supervised representation learning technique for SER. This indicates that VQ-MAE-AV learns audiovisual representations that are effectively transferable to the emotion recognition task, resulting in improved performance compared to audiovisual state-of-the-art methods.
4.4 Ablation study and model properties
We conduct a series of experiments to evaluate the impact of various hyperparameters and model designs on the emotion recognition performance of the proposed VQ-MAE-AV model. In the following paragraphs, we will present and discuss the findings of these experiments. All the ablation study is done on RAVDESS.
4.4.1 Impact of pre-training and fine-tuning
Table 7 shows the significance of pre-training and fine-tuning the VQ-MAE-AV model for audiovisual SER. Pre-training the model for the unmasking task on the VoxCeleb2 database substantially improves the emotion recognition performance, with the accuracy rising from $29.6$% to $81.5$%. Fine-tuning the encoder is also essential, as keeping it frozen leads to a $11$% drop in accuracy.
4.4.2 Impact of the encoder depth
Table 7 shows the impact of encoder depth on emotion recognition performance by varying the number of attention blocks ($L$), where a model with n attention blocks in the encoder will be denoted by VQ-MAE-AV-n. The results show that increasing the number of blocks in the encoder leads to improved performance to a certain extent. However, when the number of blocks becomes too high (as in the case of VQ-MAE-AV-20), there is no further improvement, and performance actually decreases by $1.1$% accuracy compared to VQ-MAE-AV-16.
4.4.3 Impact of the contrastive learning
The influence of contrastive learning on emotion recognition is illustrated in Table 7. When training VQ-MAE-AV exclusively with the contrastive loss, the achieved accuracy is $75.2$%, which is $9.1$% lower than when using only the generative loss. Notably, the combination of both losses, contrastive and generative, leads to improved emotion recognition, resulting in a $0.5$% accuracy boost compared to using the generative loss alone.
4.4.4 Impact of the encoder/decoder architecture
Table 7 shows the performance of the four different configurations of the VQ-MAE-AV model described in Section 3.5, including all possible combinations of the fusion strategies for the VQ-MAE-AV encoder and decoder: self-attention fusion for both the encoder and decoder (SAF-SAF); self-attention fusion for the encoder and cross-attention fusion for the decoder (SAF-CAF); cross-attention fusion for the encoder and self-attention fusion for the decoder (CAF-SAF); and cross-attention fusion for both the encoder and decoder (CAF-CAF). Among these configurations, CAF-CAF achieves the highest accuracy with a $1.5$% improvement over SAF-SAF, followed by CAF-SAF with a $1.3$% improvement over SAF-SAF, and then SAF-CAF with only a $0.3$% improvement. The cross-attention fusion encoder architecture achieves the best performance in emotion recognition, as shown by the CAF-CAF and CAF-SAF configurations. However, there exists a trade-off between performance and the number of model parameters. Notably, CAF-CAF involves slightly more than twice as many parameters as SAF-SAF.
4.4.5 Impact of the pooling strategy
Table 7 presents the impact of various pooling strategies when fine-tuning the model for audiovisual SER. The results reveal that attention-based pooling techniques, such as attention pooling and Query2Emo, outperform mean pooling. Among the two attention-based methods, Query2Emo outperforms attention pooling, with an accuracy gain of $2.8$%.
4.4.6 Impact of the audio and visual discrete token size
Fig. 5 shows the impact of the dimensions of the discrete tokens for the visual ($h$ and $w$) and audio ($d$) modalities on the audiovisual SER performance. The values of $h$ and $w$ represent the visual token size on the horizontal and vertical axes, respectively, while $d$ represents the token size on the frequency axis of the audio modality. Our study reveals that the performance of emotion recognition is impacted by both ($h,w$) and $d$ values. Therefore, it is important to select these parameters carefully. Based on our experiments, we recommend fixing them to ($h=w=4,~{}d=4$).
Figure 5: Impact of the discrete audio and visual token size on emotion recognition.
4.4.7 Impact of the audiovisual inputs
Table 7 compares our approach VQ-MAE-AV with its unimodal versions (VQ-MAE-V and VQ-MAE-A). The VQ-MAE-AV model exhibits improved performance over VQ-MAE-V and VQ-MAE-A, with gains of $5.6$% and $8.3$% in accuracy, respectively. This indicates the effectiveness of incorporating audiovisual information in the self-supervised paradigm, leading to better performance in emotion recognition.
5 Conclusion
Masked autoencoder modeling is a versatile self-supervised learning approach that can be adapted to various types of data. This paper introduced the VQ-MAE-AV model for learning representations of audiovisual speech data, which could be extended to other multimodal sequential data. VQ-MAE-AV took as input a discrete audio representation and a discrete visual representation obtained via two separate VQ-VAEs. These representations were then divided into multiple discrete tokens, with spatio-temporal tokens for the visual modality and spectro-temporal tokens for the audio modality. Pre-trained on the VoxCeleb2 dataset and fine-tuned on standard emotional audiovisual speech datasets, the experiments showed that the VQ-MAE-AV model effectively combined the audio and visual modalities for audiovisual SER, outperforming several state-of-the-art methods.
For future work, we plan to investigate other masking strategies and increase the resolution of the images to further improve emotion recognition performance. We could also explore using the representation learned by the proposed VQ-MAE-AV model for other downstream tasks in audiovisual speech processing, such as speaker recognition.
Acknowledgments
This work was performed using computational resources from the “Mésocentre” computing center of Université Paris-Saclay, CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-saclay.fr/)

### v3

@@
-
+A vector quantized masked autoencoder for audiovisual speech emotion recognition
 Abstract
-The limited availability of labeled data is a major challenge in audiovisual speech emotion recognition (SER).
-Self-supervised learning approaches have recently been proposed to mitigate the need for labeled data in various applications.
-This paper proposes the VQ-MAE-AV model, a vector quantized masked autoencoder (MAE) designed for audiovisual speech self-supervised representation learning and applied to SER.
-Unlike previous approaches, the proposed method employs a self-supervised paradigm based on discrete audio and visual speech representations learned by vector quantized variational autoencoders.
-A multimodal MAE with self- or cross-attention mechanisms is proposed to fuse the audio and visual speech modalities and to learn local and global representations of the audiovisual speech sequence, which are then used for an SER downstream task.
-Experimental results show that the proposed approach, which is pre-trained on the VoxCeleb2 database and fine-tuned on standard emotional audiovisual speech datasets, outperforms the state-of-the-art audiovisual SER methods.
-Extensive ablation experiments are also provided to assess the contribution of the different model components.
+An important challenge in emotion recognition is to develop methods that can leverage unlabeled training data.
+In this paper, we propose the VQ-MAE-AV model, a self-supervised multimodal model that leverages masked autoencoders to learn representations of audiovisual speech without labels.
+The model includes vector quantized variational autoencoders that compress raw audio and visual speech data into discrete tokens.
+The audiovisual speech tokens are used to train a multimodal masked autoencoder that consists of an encoder-decoder architecture with attention mechanisms.
+The model is designed to extract both local (i.e., at the frame level) and global (i.e., at the sequence level) representations of audiovisual speech.
+During self-supervised pre-training, the VQ-MAE-AV model is trained on a large-scale unlabeled dataset of audiovisual speech, for the task of reconstructing randomly masked audiovisual speech tokens and with a contrastive learning strategy.
+During this pre-training, the encoder learns to extract a representation of audiovisual speech that can be subsequently leveraged for emotion recognition.
+During the supervised fine-tuning stage, a small classification model is trained on top of the VQ-MAE-AV encoder for an emotion recognition task.
+The proposed approach achieves state-of-the-art emotion recognition results across several datasets in both controlled and in-the-wild conditions.
+Keywords: Self-supervised learning; Masked autoencoder; Audiovisual speech representation learning; Emotion recognition.
 keywords:
-Self-supervised learning , Masked autoencoder , Audiovisual speech processing, Emotion recognition
+Self-supervised learning , masked autoencoder , audiovisual speech representation learning , emotion recognition.
+journal: Computer Vision and Image Understanding
 […]
@@
 1 Introduction
-Emotion recognition [El Ayadi et al., 2011] aims to automatically identify emotions from unimodal or multimodal data.
-With the increasing prevalence of audiovisual interfaces, there is a growing need for systems that can accurately recognize emotions from audiovisual speech data.
-The complementarity of the audio and visual modalities has been demonstrated in various tasks [Gan et al., 2020, Gao & Grauman, 2019, Zhao et al., 2019, Sadok et al., 2024], and using them together can lead to more effective emotion recognition systems.
-According to Mehrabian [2017], only a small proportion of feelings and emotions is conveyed through the words we use in oral communication.
-The majority of our feelings and emotions is conveyed through tone, voice, facial expressions, and body language.
-In recent years, advances in artificial intelligence technology and hardware acceleration have enabled a shift towards multimodal processing in the field of emotion recognition [Ramachandram & Taylor, 2017, Tsai et al., 2019, Schoneveld et al., 2021].
-Supervised learning using large annotated datasets remains the dominating paradigm.
-However, collecting a substantial amount of emotionally labeled data can be expensive, time-consuming, and sometimes impractical.
-Most datasets for emotion recognition rely on actors to perform various emotions at specific intensities, which requires a considerable amount of acquisition time and resources.
-In addition, in-the-wild datasets suffer from annotation issues, where annotators fail to reach a consensus, resulting in a significant amount of unusable data [Busso et al., 2008].
-Unsupervised and self-supervised learning (SSL) approaches are natural ways to address these issues and have been explored in recent studies [Bengio et al., 2006, Wang et al., 2021, Dib et al., 2023],[Chen & Rudnicky, 2023].
-SSL offers the advantage of high scalability, as the SSL task can be carried out on large amounts of unlabeled speech data [Liu et al., 2022, Zhang et al., 2022].
-These methods involve pre-training models with pretext tasks and then fine-tuning them on a smaller set of labeled data for the emotion recognition task [Gong et al., 2022a, Pepino et al., 2021, Jegorova et al., 2023].
+Emotions are primarily communicated through nonverbal cues, such as tone, voice, facial expressions, and body language, rather than the words we use in oral communication (Mehrabian, 2017).
+With the increasing prevalence of audiovisual interfaces, there is a growing demand for systems capable of accurately recognizing emotions from audiovisual speech signals.
+Indeed, combining audio and visual modalities has proven to be highly effective in various tasks, in particular speech emotion recognition (SER) (El Ayadi et al., 2011; Gao and Grauman, 2019; Zhao et al., 2019; Sadok et al., 2024; Ramachandram and Taylor, 2017; Tsai et al., 2019; Schoneveld et al., 2021).
+Supervised learning using labeled datasets is the dominant machine learning paradigm in emotion recognition.
+However, obtaining datasets with emotion labels is often resource-intensive and impractical to scale.
+In fact, many emotion recognition datasets rely on actors simulating emotions, which demands significant time and effort during data collection.
+Another strategy consists of manually collecting and annotating in-the-wild data.
+However, emotion labeling is an ambiguous task, and annotators may not reach a consensus (Busso et al., 2008).
+Therefore, one important challenge is to develop methods that can leverage unlabeled training data.
+Self-supervised learning (SSL) has recently emerged as a promising technique to address this challenge (Wang et al., 2021; Dib et al., 2023; Chen and Rudnicky, 2023; Liu et al., 2022; Zhang et al., 2022).
+SSL models are pre-trained on a large-scale unlabeled dataset to solve a pretext task, and they are subsequently fine-tuned on a small amount of labeled data to solve a given task (Gong et al., 2022a; Pepino et al., 2021; Jegorova et al., 2023).
+In this paper, we present the first multimodal SSL approach based on masked autoencoders (He et al., 2022) for emotion recognition.
+We propose the VQ-MAE-AV model, a vector quantized (VQ) masked autoencoder (MAE) designed for audiovisual (AV) speech representation learning and applied to emotion recognition.
+The proposed approach first involves quantizing the audio and visual speech modalities to create discrete tokens, using vector quantized variational autoencoders (VQ-VAEs).
+Then, we pre-train the VQ-MAE-AV model in a self-supervised manner to reconstruct audiovisual speech tokens from partially visible inputs, and with an additional contrastive learning strategy.
+The VQ-MAE-AV model is based on an encoder-decoder architecture with attention mechanisms to fuse the modalities.
+Its self-supervised pre-training leverages 1 000 hours of audiovisual speech without needing emotion labels, allowing the VQ-MAE-AV model to learn an internal representation of audiovisual speech.
+Finally, this learned representation is used as input for an emotion recognition model trained on small-scale audiovisual speech datasets labeled with emotion categories.
+The experimental results demonstrate that the proposed VQ-MAE-AV model achieves superior performance compared to state-of-the-art methods across several datasets in both controlled and in-the-wild conditions.
+Additionally, extensive ablation experiments are presented to investigate the impact of different model designs.
+The paper is organized as follows.
+In Section 2 we present the related work.
+The proposed VQ-MAE-AV model is introduced in Section 3.
+Experiments are presented in Section 4 and we conclude in Section 5.
+The code and qualitative results are available at https://samsad35.github.io/VQ-MAE-AV.
 (a) Discrete visual tokens.
 (b) Discrete audio tokens.
-Figure 1: Discrete audio and visual tokens creation: (i) fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1); (ii) discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2).
-In this paper, we introduce the VQ-MAE-AV model, a vector quantized (VQ) masked autoencoder (MAE) for audiovisual (AV) speech representation learning applied to audiovisual SER.
-The overall architecture of the model is illustrated in Fig.
-1 and 2.
-To the best of our knowledge, this is the first multimodal SSL approach based on MAEs [He et al., 2022] that is proposed for audiovisual SER.
-Moreover, unlike existing multimodal MAEs proposed in other application contexts [Bachmann et al., 2022, Geng et al., 2022], the VQ-MAE-AV model operates on discrete latent representations of the modalities and introduces global tokens that are learned using contrastive learning to capture the overall sequence-wise information from each modality.
-The VQ-MAE-AV model is pre-trained on the VoxCeleb2 dataset [Chung et al., 2018] and fine-tuned on two standard emotional audiovisual speech datasets.
-The experimental results demonstrate that, across the two test datasets, the proposed VQ-MAE-AV model outperforms the state-of-the-art audiovisual SER methods.
-Extensive ablation experiments are also presented to study the impact of different model designs.
-The code and qualitative results are available at https://samsad35.github.io/VQ-MAE-AV.
-Figure 2: VQ-MAE-AV model structure. See the first paragraph of Section 3 for a complete description of the pipeline.
+Fig.
+1: Discrete audio and visual tokens creation: (i) fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1); (ii) discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2).
 2 Related work
 Self-supervised learning approaches
-SSL models can be broadly categorized into discriminative and generative approaches [Zhang et al., 2022].
-Discriminative SSL focuses on creating pairs or groups of data samples and formulating loss functions that enable the model to distinguish or group these samples, which can later benefit downstream tasks [Chen et al., 2020b, a].
-Pretext tasks can consist of solving jigsaw puzzles [Noroozi & Favaro, 2016] or predicting image rotations Gidaris et al.
-[2018] for example.
-Contrastive learning has emerged as the predominant paradigm in discriminative SSL [Alayrac et al., 2020, Akbari et al., 2021].
-In contrast, generative SSL involves generating or reconstructing segments of unlabeled and potentially corrupted data using an autoencoder model [He et al., 2022, Bao et al., 2021, Xie et al., 2022].
-The latent representation produced by the encoder can subsequently be used for downstream tasks.
+SSL models can be broadly classified into two main categories: discriminative and generative approaches (Zhang et al., 2022).
+Discriminative SSL focuses on creating pairs or groups of data samples and formulating loss functions that allow the model to differentiate or group these samples, which can later benefit downstream tasks (Chen et al., 2020b, a).
+For instance, pretext tasks can consist of solving jigsaw puzzles (Noroozi and Favaro, 2016) or predicting image rotations (Gidaris et al., 2018).
+Contrastive learning has emerged as the predominant paradigm in discriminative SSL (Alayrac et al., 2020; Akbari et al., 2021).
+In contrast, generative SSL involves generating or reconstructing segments of unlabeled and potentially corrupted data using an encoder-decoder model (He et al., 2022; Bao et al., 2021; Xie et al., 2022).
+The latent representation produced by the encoder can then be leveraged for downstream tasks.
 Masked autoencoder
-The present paper focuses on the MAE, an SSL generative model that uses an asymmetric encoder-decoder architecture with input masking [He et al., 2022].
-The MAE approach is inspired by masked language modeling [Devlin et al., 2019] and has been successfully applied to image modeling thanks to the development of Vision Transformers (ViT) [Dosovitskiy et al., 2020].
-The MAE has recently been adapted for audio using a 2D time-frequency representation [Gong et al., 2022a, Baade et al., 2022, Huang et al., 2022].
+The present paper focuses on the MAE, an SSL generative model that employs an asymmetric encoder-decoder architecture with input masking (He et al., 2022).
+The MAE approach is inspired by the concept of masked language modeling (Devlin et al., 2019) and has been successfully applied to image modeling thanks to the development of the Vision Transformer (ViT) (Dosovitskiy et al., 2020).
+The MAE has recently been extended to audio using a 2D time-frequency representation (Gong et al., 2022a; Baade et al., 2022; Huang et al., 2022).
 In the MAE paradigm, the input is divided into non-overlapping patches, each represented by a token embedding.
-A proportion of tokens are masked, typically 75% for image/audio modeling and 15% for text modeling, and only the visible tokens are fed to the encoder.
+Some tokens are masked, typically 75% for image/audio modeling and 15% for text modeling, and only the visible tokens are fed to the encoder.
 The encoder outputs are used to reconstruct the masked tokens through a lightweight decoder.
-To successfully reconstruct the masked tokens, the encoder must generate a semantic representation, which can then be effectively transferred to downstream tasks [He et al., 2022].
-It was recently shown that combining the task of reconstructing masked tokens with contrastive learning can improve the representation learned by an MAE [Huang et al., 2023, Gong et al., 2022a].
-Extensions of MAE for sequential and multimodal representation learning
-A recent extension of the MAE was presented in [Feichtenhofer et al., 2022, Tong et al., 2022] for modeling image sequences, called Video-MAE.
-It uses the same architecture as the vanilla MAE [He et al., 2022] but incorporates a masking process from video ViT (ViViT) [Arnab et al., 2021].
-Since videos often contain redundant information, particularly in scenes with no motion, the authors propose a cubic masking approach along the temporal axis, combined with a high masking ratio of 90%.
-Other works have extended the MAE to handle multimodal data [Bachmann et al., 2022, Geng et al., 2022, Gong et al., 2022b].
-MultiMAE [Bachmann et al., 2022] encodes a small random subset of visible tokens from multiple modalities (RGB, depth, and semantic images) and is trained to reconstruct the missing ones.
-M3AE [Geng et al., 2022] is a unified MAE architecture for two input modalities (image and text).
-The main difference in architecture between M3AE and MultiMAE lies in the architecture of the decoder.
-The MulitMAE approach introduces individual decoders for each modality, enhancing their fusion by incorporating a cross-attention layer at the outset of each decoder.
+To successfully reconstruct the masked tokens, the encoder must capture a semantic representation, which can then be effectively transferred to downstream tasks (He et al., 2022).
+It was recently shown that combining the task of reconstructing masked tokens with contrastive learning can improve the representation learned by an MAE (Huang et al., 2023; Gong et al., 2022a).
+Extensions of the MAE for sequential and multimodal representation learning
+A recent extension of the MAE was presented in (Feichtenhofer et al., 2022; Tong et al., 2022) for modeling image sequences, called Video-MAE.
+This method employs the same architecture as the vanilla MAE (He et al., 2022) but incorporates a masking process from video ViT (ViViT) (Arnab et al., 2021).
+Since videos often contain redundant information, particularly in scenes with no motion, the authors proposed a cubic masking approach along the temporal dimension, combined with a high masking ratio of 90%.
+Other works have extended the MAE to handle multimodal data (Bachmann et al., 2022; Geng et al., 2022; Gong et al., 2022b).
+MultiMAE (Bachmann et al., 2022) encodes a small random subset of visible tokens from multiple modalities (RGB, depth, and semantic images) and trains the model to reconstruct the missing tokens.
+M3AE (Geng et al., 2022) is a unified MAE architecture for two input modalities (image and text).
+The main difference between M3AE and MultiMAE lies in the architecture of the decoder.
+The MulitMAE approach introduces individual decoders for each modality, enhancing their fusion by incorporating a cross-attention layer at the beginning of each decoder.
 On the other hand, the M3AE approach uses a single decoder that takes the concatenated tokens from all modalities as input.
@@
 In the literature, MAEs are typically trained using the L1 or L2 losses, which can negatively affect the reconstruction quality of the masked tokens, resulting, for instance, in blurred or noisy images or sounds.
-It has been shown that improving the quality of MAE reconstructions can be beneficial in terms of downstream task performance [He et al., 2022].
-Several approaches have been proposed in that sense, which include adding a perceptual loss [Dong et al., 2021] or using discrete representations obtained from VQ generative adversarial networks (VQ-GANs) [Esser et al., 2021] or variational autoencoders (VQ-VAEs) [Van Den Oord et al., 2017] to train the MAE [Li et al., 2023, Sadok et al., 2023].
-These works only considered an unimodal setting, while the present paper proposes a multimodal MAE for audiovisual speech representation learning.
-Recently, a model called CAV-MAE [Gong et al., 2022b] was proposed combining MAEs and contrastive learning to learn a representation from audiovisual data.
-However, different from the proposed VQ-MAE-AV model, CAV-MAE processes raw audiovisual data instead of vector-quantized latent representations, it uses simple mean pooling strategies instead of attention pooling, and it is not applied to SER.
+Studies have demonstrated that improving the quality of MAE reconstructions can be beneficial in terms of downstream task performance (He et al., 2022).
+Several approaches have been proposed for that purpose, for instance adding a perceptual loss (Dong et al., 2021) or using discrete representations obtained from VQ generative adversarial networks (VQ-GANs) (Esser et al., 2021) or variational autoencoders (VQ-VAEs) (Van Den Oord et al., 2017) to train the MAE (Li et al., 2023; Sadok et al., 2023).
+While these works only considered a unimodal setting, the present paper proposes a multimodal MAE for audiovisual speech representation learning.
+Recently, a model called CAV-MAE (Gong et al., 2022b) was proposed combining MAEs and contrastive learning to learn a representation from audiovisual data.
+However, unlike the proposed VQ-MAE-AV model, CAV-MAE processes raw audiovisual data instead of vector-quantized latent representations, employs simple mean pooling strategies rather than attention pooling, and has not been applied to SER.
+Fig. 2: VQ-MAE-AV model structure. See the first paragraph of Section 3 for a complete description of the pipeline.
 3 The VQ-MAE-AV model
-This section presents the VQ-MAE-AV model. The overall approach is illustrated in Fig. 1 and 2 and it can be summarized as follows:
+This section presents the VQ-MAE-AV model, which is illustrated in Fig. 1 and 2 and can be summarized as follows:
   • 1.
@@
   • 8.
-    After self-supervised learning, the VQ-MAE-AV encoder and attention pooling layers are fine-tuned for supervised audiovisual SER (see Section 3.7).
+    After self-supervised learning, a small classification model is trained and the VQ-MAE-AV encoder is fine-tuned for supervised audiovisual SER (see Section 3.7).
 This section will present each above-listed aspect of the model in more detail.
 3.1 Vector quantized variational autoencoder
-The proposed multimodal self-supervised approach uses the discrete latent representation of two pre-trained and frozen VQ-VAEs [Van Den Oord et al., 2017].
+The proposed multimodal self-supervised approach uses the discrete latent representation of two pre-trained and frozen VQ-VAEs (Van Den Oord et al., 2017).
 Specifically, as illustrated in Fig.
-1, we use the VQ-VAE-audio and VQ-VAE-visual encoders to obtain compressed and quantized representations of the input speech power spectrogram ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}\in\mathbb{R}^{T_{a}\times D}$ where $T_{a}$ and $D$ correspond to the time and frequency dimensions, and of the input image sequence ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}\in\mathbb{R}^{T_{v}\times H\times W\times C}$ where $T_{v}$, $H$, $W$ and $C$ correspond to the time, height, width, and channel dimensions.
-The audio and visual quantized representations are denoted by ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$, respectively.
+1, we use the VQ-VAE-A (A for audio) and VQ-VAE-V (V for visual) encoders to obtain compressed and quantized representations of the input speech power spectrogram ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}\in\mathbb{R}^{T_{a}\times D}$ and of the input image sequence ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}\in\mathbb{R}^{T_{v}\times H\times W\times C}$, where $T_{a}$ and $D$ correspond to the time and frequency dimensions of the audio modality, and $T_{v}$, $H$, $W$ and $C$ correspond to the time, height, width, and channel dimensions of the audio modality.
+The audio and visual quantized representations are denoted by ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{T_{v}\times H^{\prime}\times W^{\prime}}$, respectively.
 Each entry of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ corresponds to the index of a vector in the VQ-VAE codebooks.
 Notably, ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ retains the time-frequency structure of the original spectrogram, while ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ retains the spatio-temporal structure of the original sequence images.
-This is because the VQ-VAE-audio and VQ-VAE-visual models are designed to be fully convolutional on the frequency and spatial axes, respectively, and they process the frames within a sequence independently.
+This is because the VQ-VAE-A and VQ-VAE-V models are designed to be fully convolutional on the frequency and spatial axes, respectively, and they process the frames within a sequence independently.
 Therefore, compression occurs along the frequency axis ($D^{\prime}\ll D$) for ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and along the x and y-axes of the image ($H^{\prime}\ll H$, $W^{\prime}\ll W$) for ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$.
@@
 The training procedure of the VQ-VAEs follows the original approach presented in Van Den Oord et al.
-[2017].
-In particular, the VQ-VAE loss functions involve a reconstruction term between the original and reconstructed data, which corresponds to the mean squared error for the visual modality and to the Itakura-Saito divergence for the audio modality [Févotte et al., 2009].
+(2017).
+In particular, the VQ-VAE loss functions involve a reconstruction term between the original and reconstructed data, which corresponds to the mean squared error for the visual modality and to the Itakura-Saito divergence for the audio modality (Févotte et al., 2009).
 More details are provided in Section 4.1.2.
@@
 As shown in Fig.
-1, the audio and visual quantized representations ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$ from the output of the VQ-VAE encoders are divided into non-overlapping patches to build discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot t_{a})\times(n_{d}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot t_{v})\times(n_{h}\cdot h)\times(n_{w}\cdot w)}$, where $T_{a}=n_{t_{a}}\cdot t_{a}$, $D^{\prime}=n_{d}\cdot d$, $T_{v}=n_{t_{v}}\cdot t_{v}$, $H^{\prime}=n_{h}\cdot h$, and $W^{\prime}=n_{w}\cdot w$.
-These representations are reshaped to ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$, which are seen as sequences of $n_{t_{a}}\cdot n_{d}$ and $n_{t_{v}}\cdot n_{h}\cdot n_{w}$ tokens of dimension $t_{a}\cdot d$ and $t_{v}\cdot h\cdot w$, respectively.
+1, the audio and visual quantized representations ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{T_{v}\times H^{\prime}\times W^{\prime}}$ from the output of the VQ-VAE encoders are divided into non-overlapping patches to build discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot t_{a})\times(n_{d}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{(n_{t_{v}}\cdot t_{v})\times(n_{h}\cdot h)\times(n_{w}\cdot w)}$, where $T_{a}=n_{t_{a}}\cdot t_{a}$, $D^{\prime}=n_{d}\cdot d$, $T_{v}=n_{t_{v}}\cdot t_{v}$, $H^{\prime}=n_{h}\cdot h$, and $W^{\prime}=n_{w}\cdot w$.
+These representations are reshaped to ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$, which are seen as sequences of $n_{t_{a}}\cdot n_{d}$ and $n_{t_{v}}\cdot n_{h}\cdot n_{w}$ tokens of dimension $t_{a}\cdot d$ and $t_{v}\cdot h\cdot w$, respectively.
 The use of discrete audio and visual tokens in VQ-MAE-AV has several motivations.
 Firstly, by dividing the audiovisual data into spatio-spectro-temporal patches, the method could learn to relate audio tokens to visual tokens.
-For example, the audio tokens are expected to correlate strongly with the visual tokens corresponding to the mouth area [Arnela et al., 2016, Sadok et al., 2024].
+For example, the audio tokens are expected to correlate strongly with the visual tokens corresponding to the mouth area (Arnela et al., 2016; Sadok et al., 2024).
 The proposed method can potentially learn a representation that captures shared and distinctive information between the two modalities by exploiting their complementarity.
-Additionally, the use of discrete tokens can reduce the computational cost of the method as it involves working with a reduced representation of the data, which allows us to increase the number of tokens (i.e., manipulate longer audiovisual speech sequences) without exploding in the number of trainable parameters compared to the multimodal MAE in the literature [Bachmann et al., 2022].
+Additionally, the use of discrete tokens can reduce the computational cost of the method as it involves working with a reduced representation of the data, which allows us to increase the number of tokens (i.e., manipulate longer audiovisual speech sequences) without exploding in the number of trainable parameters compared to the multimodal MAE in the literature (Bachmann et al., 2022).
 3.3 Masking
-We apply masking to the audio and visual sequences of tokens.
-The MAE aims to reconstruct the masked tokens from the visible tokens to learn a semantic representation of the data.
-The masking strategy impacts the performance of downstream tasks.
-In the original MAE, the masked tokens are randomly drawn with a target ratio of the entire set of tokens, typically 75%.
-In multimodal scenarios, the basic strategy would be to do the same for each modality, but it has been shown that it is more interesting to implement a coupled masking strategy between the modalities [Bachmann et al., 2022].
-This involves randomly drawing a masking proportion $p_{a}\in[0,1]$ for the audio modality and $p_{v}\in[0,1]$ for the visual modality such that $p_{a}+p_{v}=1$.
-To do this, a Dirichlet distribution is used: $(p_{a},p_{v})\sim Dir(\alpha_{a},\alpha_{v})$ where $\alpha_{a},\alpha_{v}>0$ are the concentration parameters of the distribution.
-When $\alpha_{a}=\alpha_{v}=1$, the distribution is uniform overall points in its support (the 1-simplex), and $\alpha_{a}\gg\alpha_{b}$ results in a sampling behavior where most of the tokens are taken from the audio modality, and vice versa.
-These parameters can be used to reduce the dominance of one modality over the other.
-This strategy allows the reconstruction of missing information from one modality by relying on another.
-Therefore, the model will leverage the less masked modality for the reconstruction of the more masked one.
-For this paper, we set $\alpha_{a}=\alpha_{v}=1$.
+The masking strategy employed to train the MAE will impact the performance of downstream tasks He et al.
+(2022).
+In the original MAE, the masked tokens are chosen randomly given a target masking ratio, typically 75%.
+The MAE is then trained to reconstruct the masked tokens from the visible ones, and in doing so it learns a representation that can be used for downstream tasks.
+In a multimodal scenario, the straightforward approach would be to apply this same masking strategy independently on each modality.
+However, it has been shown that it is more effective to implement a coupled masking strategy between the modalities (Bachmann et al., 2022).
+Therefore, to the train the VQ-MAE-AV model, the masking ratio for the audio and visual modalities is drawn randomly according to a uniform distribution on the 1-simplex.
+This means that if $p\times 100~{}\%$ of the tokens are masked in one modality, then $(1-p)\times 100~{}\%$ of the tokens are masked in the other modality, where $p$ is distributed uniformly between 0 and 1.
+This strategy allows the reconstruction of missing information from one modality by relying on the other.
+Fig. 3: Overview of the three emotion recognition models trained on top of the VQ-MAE-AV encoder.
 3.4 Continuous embedding vectors
 The discrete tokens correspond to the indices obtained through the quantization step of the pretrained VQ-VAE encoder.
-Before being input to the VQ-MAE-AV encoder, these discrete tokens are replaced with trainable continuous embedding vectors taken from an audio codebook in $\mathbb{R}^{k_{a}\times e_{a}}$ and from a visual codebook in $\mathbb{R}^{k_{v}\times e_{v}}$, where $k_{a/v}$ is the number of codes in the codebook and $e_{a/v}$ is the dimension of each code.
+Before being input into the VQ-MAE-AV encoder, these discrete tokens are replaced with trainable continuous embedding vectors taken from an audio codebook in $\mathbb{R}^{k_{a}\times e_{a}}$ and from a visual codebook in $\mathbb{R}^{k_{v}\times e_{v}}$, where $k_{a/v}$ is the number of codes in the codebook and $e_{a/v}$ is the dimension of each code.
 This is simply achieved by replacing the indices of a discrete token with the corresponding vectors of dimension $e_{a/v}$ in the codebook.
-After this embedding process, the sequences of discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ in $\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$ are transformed into sequences of continuous tokens ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times({t_{a}}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$.
+After this embedding process, the sequences of discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ in $\mathbb{N}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$ are transformed into sequences of continuous tokens ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times({t_{a}}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{R}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$.
+The dimensions $e_{a}$ and $e_{v}$ of the audio and visual codes are chosen such that the continuous tokens in both modalities have the same dimension $E={t_{a}}\cdot d\cdot e_{a}={t_{v}}\cdot h\cdot w\cdot e_{v}$, which is necessary for concatenation in the VQ-MAE-AV encoder (see the next subsection).
+Therefore, we have ${\color[rgb]{1,0,0}\mathbf{x}_{c}^{(a)}}\in\mathbb{R}^{(n_{t_{a}}\cdot n_{d})\times E}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{c}^{(v)}}\in\mathbb{R}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times E}$.
 3.5 VQ-MAE-AV encoder and decoder
 3.5.1 Attention Block
-The VQ-MAE-AV encoder and decoder are built with multi-head Attention blocks similar to those used in the Vision Transformer (ViT) [Dosovitskiy et al., 2020].
+The VQ-MAE-AV encoder and decoder are built with multi-head Attention blocks similar to those used in the Vision Transformer (ViT) (Dosovitskiy et al., 2020).
 Each block comprises a multi-head attention layer, normalization layers, and a Multi-layer Perceptron (MLP).
@@
 2, and they will be used to capture the inter and intra-relationships between audio and visual tokens.
-This attention block is inspired by the attention layer in the original transformer [Vaswani et al., 2017].
+This attention block is inspired by the attention layer in the original transformer (Vaswani et al., 2017).
 To simplify the reading afterward, we denote the attention block by $\texttt{Attention}(Q,V,K)$, where $Q,V,K$ are the query, value, and key, respectively.
@@
 We propose two fusion strategies of the audio and visual speech data, resulting in two architectures for the VQ-MAE-AV encoder.
-The first fusion strategy is called self-attention fusion (SAF).
+The first fusion strategy is called self-attention fusion.
 As represented in Fig.
@@
 This concatenation operates on the first dimension of the variables, i.e., we obtain a sequence of $(n_{t_{a}}\cdot n_{d})+(n_{t_{v}}\cdot n_{h}\cdot n_{w})$ tokens after concatenation.
-The second fusion strategy is called cross-attention fusion (CAF).
+The second fusion strategy is called cross-attention fusion.
 As shown in Fig.
@@
 3.5.3 Global tokens
-In addition to the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, we learn two sequence-wise global tokens, denoted by ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$, which can be thought of as similar to [CLS] tokens [He et al., 2022].
+In addition to the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, we learn two sequence-wise global tokens, denoted by ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}\in\mathbb{R}^{1\times E}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}\in\mathbb{R}^{1\times E}$, which can be thought of as similar to [CLS] tokens (He et al., 2022).
 These global modality-specific tokens are introduced to aggregate the spectro-temporal and spatio-temporal information in the two modalities, which can be useful for downstream tasks involving predictions at the sequence level, such as audiovisual SER.
@@
 ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}$ $=\texttt{Attention}\left(Q_{(v)},V_{(v)},K_{(v)}\right),$ (2)
-where $Q_{(a)}=\mathcal{Z}^{(a)}\in\mathbb{R}^{1\times(t_{a}\cdot d\cdot e_{a})}$ and $Q_{(v)}=\mathcal{Z}^{(v)}\in\mathbb{R}^{1\times({t_{v}}\cdot h\cdot w\cdot e_{v})}$ represent respectively trainable audio and visual tokens, as proposed in Touvron et al.
-[2021], $V_{(a)}=K_{(a)}={\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$, and $V_{(v)}=K_{(v)}={\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
+where $Q_{(a)}\in\mathbb{R}^{1\times E}$ and $Q_{(v)}\in\mathbb{R}^{1\times E}$ represent respectively trainable audio and visual tokens, as proposed in Touvron et al.
+(2021), $V_{(a)}=K_{(a)}={\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$, and $V_{(v)}=K_{(v)}={\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$.
 3.5.4 Decoders
@@
 2.
-The mask tokens actually correspond to one single trainable vector as proposed in the original MAE [He et al., 2022].
+The mask tokens actually correspond to one single trainable vector as proposed in the original MAE (He et al., 2022).
 Similarly as for the encoder, the audio and visual inputs of the VQ-MAE-AV decoder can be fused using either self-attention fusion or cross-attention fusion.
@@
 The output of this linear layer corresponds to the logits of the discrete tokens.
-After applying a softmax operation, we obtain reconstructions ${\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}$ and ${\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}$ of the indices ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ that were provided by the VQ-VAE-audio and VQ-VAE-visual encoders, respectively.
-3.6 VQ-MAE-AV loss functions
-3.6.1 Generative loss function
-To train the VQ-MAE-AV model, we minimize the cross-entropy loss applied only to the masked discrete tokens:
-$\mathcal{L}_{rec}=\texttt{cross-entropy}\Biggl{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)},~{}{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}\Big{(}\Omega_{\mathcal{M}}^{(a)}\Big{)}\Biggl{)}$
-[…]
-where $\mathbf{x}\left(\Omega_{\mathcal{M}}^{(\cdot)}\right)$ denotes the set of masked tokens in $\mathbf{x}$.
-Another benefit of manipulating discrete representations for multimodal inputs is the homogeneity of the losses, which does not require balancing the losses between the two modalities.
-3.6.2 Contrastive loss function
+After applying a argmax operation, we obtain reconstructions ${\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}$ and ${\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}$ of the indices ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ that were provided by the VQ-VAE-A and VQ-VAE-V encoders, respectively.
+3.6 Loss functions
+The VQ-MAE-AV model is trained in a self-supervised manner to minimize a generative loss $\mathcal{L}_{gen}$ between the reconstructed and original tokens and a contrastive loss $\mathcal{L}_{NCE}$ between the audio and visual global tokens.
+Due to their discrete nature, the cross-entropy (CE) is naturally used to measure the reconstruction quality of the audiovisual masked tokens:
+$\mathcal{L}_{gen}$ $\Big{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}},{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}},{\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}},{\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}},\Omega^{(a)},\Omega^{(v)}\Big{)}=$
+$\texttt{CE}\Biggl{(}{\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\Big{(}\Omega^{(a)}\Big{)},~{}{\color[rgb]{1,0,0}\mathbf{\hat{x}}^{(a)}_{q}}\Big{(}\Omega^{(a)}\Big{)}\Biggl{)}\,+\,\texttt{CE}\Biggl{(}{\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\Big{(}\Omega^{(v)}\Big{)},~{}{\color[rgb]{0,0,1}\mathbf{\hat{x}}^{(v)}_{q}}\Big{(}\Omega^{(v)}\Big{)}\Biggl{)},$ (3)
+where $\mathbf{x}\left(\Omega^{(\cdot)}\right)$ denotes the set of masked tokens in $\mathbf{x}$.
+As can be seen, another benefit of manipulating discrete representations for multimodal inputs is the homogeneity of the losses, which does not require balancing the losses between the two modalities.
 Building upon the approaches presented in Alayrac et al.
-[2020], Akbari et al.
-[2021], the global tokens are learned using noise contrastive estimation.
+(2020); Akbari et al.
+(2021), the global tokens are learned using noise contrastive estimation.
 This approach enhances the alignment of audiovisual speech pairs by grouping together embeddings that belong to the same time sequence and separating them from those that do not correspond to the same sequence.
-This approach involves minimizing the loss function $\mathcal{L}_{NCE}({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$, which is defined by:
-$\mathcal{L}_{NCE}(\mathbf{u},\mathbf{v})=$
-$-\log\left(\frac{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}}{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}+\sum\limits_{(\mathbf{u}^{\prime},~{}\mathbf{v}^{\prime})\in\mathcal{N}}\exp\Bigl{(}\mathbf{u}^{\prime\top}\mathbf{v}^{\prime}/\tau\Bigl{)}}\right).$ (4)
+It involves minimizing the loss function $\mathcal{L}_{NCE}({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$, which is defined by:
+$\mathcal{L}_{NCE}(\mathbf{u},\mathbf{v})=-\log\left(\frac{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}}{\exp\Bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\Bigl{)}+\sum\limits_{(\mathbf{u}^{\prime},~{}\mathbf{v}^{\prime})\in\mathcal{N}}\exp\Bigl{(}\mathbf{u}^{\prime\top}\mathbf{v}^{\prime}/\tau\Bigl{)}}\right).$ (4)
 To form positive pairs $({\color[rgb]{1,0,0}\mathbf{w}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{w}^{(v)}})$ for both audio and visual modalities, we select corresponding streams from the same temporal location in the video.
@@
 The sensitivity of the NCE loss in distinguishing between positive and negative pairs is regulated by a temperature parameter $\tau\in\mathbb{R}$.
-3.6.3 Final loss function
-The final loss function corresponds to the sum of the generative loss function in equation (3.6.1) and of the contrastive loss function in equation (3.6.2).
-(a) Qualitative unmasking results for the audio modality.
+The overall loss function used to train VQ-MAE-AV is simply the sum of the generative and contrastive loss functions in (3.6) and (4).
+3.7 Emotion recognition
+After self-supervised learning on a large-scale unlabeled dataset, the VQ-MAE-AV model is used for emotion recognition.
+The task is to predict the emotion class of an input audiovisual speech sequence.
+To address this task, we need to introduce a small emotion recognition model for predicting the emotion category from the audiovisual speech representation provided by the pre-trained VQ-MAE-AV encoder.
+In this work, we propose and investigate three different emotion recognition models, which are described in the remainder of this section and illustrated in Fig.
+3.
+All emotion recognition models are trained for a supervised classification task using the asymmetric loss of Ridnik et al.
+(2021).
+Along with the training of the emotion recognition model, we also fine-tune the VQ-MAE-AV encoder.
+3.7.1 Emotion recognition model based on attention pooling
+The first approach to performing emotion recognition from the proposed VQ-MAE-AV model is to use the sequence-wise global tokens ${\color[rgb]{1,0,0}\mathbf{w}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{w}^{(v)}}$, which were computed using attention pooling of the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, as described in Section 3.5.3.
+In this approach, the two global tokens are concatenated and passed through a single linear layer followed by a softmax operation to predict the emotion class probabilities.
+3.7.2 Emotion recognition model based on mean pooling
+The second approach for emotion recognition replaces the previous attention pooling strategy with a more naive one, where the token-wise representations ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$ are aggregated temporally using a simple parameter-free mean pooling operation.
+The resulting vectors are then concatenated and passed through a single linear layer, as before.
+3.7.3 Emotion recognition model based on Query2Emo
+Finally, we propose a third emotion recognition approach, referred to as Query2Emo and inspired by (Liu et al., 2021).
+As illustrated in Fig.
+3, Query2Emo involves cross-attention between all audio and visual tokens (concatenation of ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) as key and value, and the emotion classes represented by trainable embeddings as the query:
+$\mathbf{w}_{emo}=\texttt{Attention}\left(Q_{emo},\mathbf{z},\mathbf{z}\right),$ (5)
+where $\mathbf{z}$ denotes the concatenation of the two sequences ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$, and $Q_{emo}\in\mathbb{R}^{K_{emo}\times E}$ corresponds to $K_{emo}$ trainable tokens of dimension $E$, with $K_{emo}$ the number of emotion classes.
+This cross-attention aims to learn the relationship between the emotion representation $Q_{emo}$ and the audiovisual tokens for the SER task.
+Query2Emo consists of two attention blocks as illustrated in Fig.
+3.
+The model provides at the output $K_{emo}$ vectors of dimension $E$ (i.e., $\mathbf{w}_{emo}\in\mathbb{R}^{K_{emo}\times E})$, which are then concatenated and passed through a single linear layer, as for the two previous emotion recognition models.
+4 Experiments
+This section starts by presenting the experimental setup, including the datasets and preprocessing, the model architecture, and the training configuration.
+Then, we present a first experiment that aims to measure the reconstruction quality of the VQ-MAE-AV model when applied to masked audiovisual speech data.
+We then evaluate the performance of the proposed approach for emotion recognition using four audiovisual speech datasets and comparing with several state-of-the-art methods.
+Finally, we present the results of an ablation study that measures the impact of various hyperparameters and architecture choices on the performance of our method.
+4.1 Experimental setup
+4.1.1 Datasets and preprocessing
+Dataset for self-supervised training
+To pre-train VQ-MAE-AV, we use the VoxCeleb2 dataset (Chung et al., 2018), which offers a broad range of audiovisual speech data from open-source media, with each video featuring a single speaker.
+We restricted our dataset use to a subset of around 1000 hours of audiovisual speech, encompassing 2170 different speakers.
+The test set includes about 100 hours of audiovisual speech data with 117 different speakers.Wild2
+Data pre-processing
+The VQ-VAE-A and VQ-VAE-V models are trained on the VoxCeleb2 dataset.
+The former is trained on short-time Fourier transform (STFT) power spectrograms (${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$), while the latter is trained on sequences of RGB images (${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$) captured at 25 fps, cropped and resized to a resolution of $96\times 96$, and aligned using Face-Alignment (Bulat and Tzimiropoulos, 2017).
+To compute the STFT, a Hann window of 64 ms (1024 samples at 16 kHz) and a $68$% overlap are used, resulting in a sample rate of 50 Hz, which is twice the sample rate of the visual modality.
+This leads to sequences of $D=513$ Fourier coefficients.
+Emotional audiovisual speech datasets
+We fine-tune and evaluate the proposed approach for audiovisual SER in both controlled and in-the-wild conditions, using four datasets that are presented below.
+  • 1.
+    RAVDESS (Livingstone and Russo, 2018) is a dataset that consists of 1,440 videos recorded by 24 English-speaking actors in controlled conditions.
+It is labeled with eight emotions: neutral, calm, happy, sad, angry, fearful, disgust, surprised.
+  • 2.
+    CREMA-D (Cao et al., 2014) is a dataset that consists of 7,442 videos recorded by 91 English-speaking actors in controlled conditions.
+Actors spoke from a selection of 12 sentences, using one of six emotions (anger, disgust, fear, happy, neutral, and sad) with four different intensities (low, medium, high, and unspecified).
+  • 3.
+    DFEW (Jiang et al., 2020) is a dataset containing 16,372 video clips of English-speaking subjects, annotated with seven emotions: neutral, happy, sad, surprise, fear, anger, and disgust.
+The clips are extracted from diverse movie scenes, ensuring a wide range of variations in expressions, lighting, and occlusions, providing a challenging benchmark for emotion recognition in the wild.
+  • 4.
+    Aff-Wild2 (Kollias and Zafeiriou, 2018): Aff-Wild2 comprises 564 videos extracted from Youtube (around 2.8M frames), including 554 different English-speaking subjects, and annotated with seven emotions: neutral, anger, disgust, fear, happiness, sadness, surprise.
+Similar to DFEW, Aff-Wild2 captures spontaneous expressions in uncontrolled environments, including diverse recording conditions, making it suitable for emotion recognition in real-world scenarios.
+We selected these datasets as they are commonly used for the emotion recognition task, and the raw audiovisual speech data is accessible.
+We performed 6-fold and 10-fold cross-validation for the RAVDESS and CREAM-D datasets to ensure a fair comparison with previous works.
+For DFEW and Aff-Wild2, the train, validation, and test splits are already defined.
+For all datasets, the speaker identities are different in the training and evaluation sets, allowing us to evaluate SER in a speaker-independent setting.
+(a) Qualitative reconstruction results for the audio modality.
 The spectrogram highlighted in the red box represents the original spectrogram.
@@
 The reconstructions using VQ-MAE-AV can be seen directly below these masked spectrograms.
-(b) Qualitative unmasking results for the visual modality.
+(b) Qualitative reconstruction results for the visual modality.
 The first sequence shows the original video, followed by the next two sequences representing the masked video with a ratio of 50% and its reconstruction using VQ-MAE-AV.
 The last two sequences represent the masked video with a ratio of 80% and its reconstruction using VQ-MAE-AV.
-Figure 3: Quantitative results of the audio reconstruction (a) and visual reconstruction (b) using the VQ-MAE-AV model.
-3.7 Fine-tuning for audiovisual SER
-After pre-training VQ-MAE-AV on token unmasking, the model is fine-tuned for audiovisual SER.
-We propose three different approaches for fine-tuning the model.
-The first two approaches rely on a pooling operation (of the encoder outputs ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) to extract the global audio and visual tokens, which is either a simple mean pooling or the attention pooling discussed in Section 3.5.3.
-The extracted tokens are then concatenated and passed through a single linear layer, followed by an $argmax$ operation to classify the emotion.
-The third proposed approach, referred to as Query2Emo and inspired by [Liu et al., 2021], involves cross-attention between all audio and visual tokens (concatenation of ${\color[rgb]{1,0,0}\mathbf{z}^{(a)}},~{}{\color[rgb]{0,0,1}\mathbf{z}^{(v)}}$) as key and value, and the emotion classes represented by trainable embeddings as the query.
-Query2Emo has a single attention block for both the encoder and decoder.
-The extracted embeddings are then concatenated and passed through a single linear layer.
-We adopt the asymmetric loss [Ridnik et al., 2021] for all these approaches between the predicted emotion and the ground-truth emotion.
-4 Experiments
-This section presents our study’s experimental setup and results, as well as a discussion of their implications.
-We evaluate the effectiveness of the proposed approach for emotion recognition on two standard audiovisual speech databases and compare it with state-of-the-art methods.
-Additionally, we perform an ablation study to analyze the impact of various hyperparameters and architectures on the performance of our method.
-We present qualitative audio and visual reconstruction results in Fig.
-3(a) and Fig.
-3(b), respectively.
-The reconstructions are illustrated for two distinct masking ratios: 50% and 80%.
-Additional qualitative results are available online (see the webpage address at the end of Section 1).
-Furthermore, we conducted a study to measure the reconstruction quality at different masking ratios of the audiovisual speech data.
-This analysis is detailed in Section 4.2 and demonstrates the effectiveness of leveraging multimodality to enhance reconstruction quality.
-4.1 Experimental setup
-4.1.1 Datasets and preprocessing
-Dataset for self-supervised training
-To pre-train VQ-MAE-AV, we use the VoxCeleb2 dataset [Chung et al., 2018], which offers a broad range of audiovisual speech data from open-source media, with each video featuring a single speaker.
-We restricted our dataset use to a subset of around 1000 hours of audiovisual speech, encompassing 2170 different speakers.
-The test set includes about 100 hours of audiovisual speech data with 117 different speakers.
-Data pre-processing
-The VQ-VAE-audio and VQ-VAE-visual models are trained on the VoxCeleb2 dataset.
-The former is trained on short-time Fourier transform (STFT) power spectrograms (${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}$), while the latter is trained on sequences of RGB images (${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}$) captured at 25 fps, cropped and resized to a resolution of $96\times 96$, and aligned using Face-Alignment [Bulat & Tzimiropoulos, 2017].
-To compute the STFT, a Hann window of 64 ms (1024 samples at 16 kHz) and a $68$% overlap are used, resulting in a sample rate of 50 Hz, which is twice the sample rate of the visual modality.
-This leads to sequences of $D=513$ Fourier coefficients.
-Emotional audiovisual speech databases
-We fine-tune and evaluate the proposed approaches on two emotional audiovisual speech databases.
-  • 1.
-    RAVDESS [Livingstone & Russo, 2018]: This English database consists of 1440 videos recorded by 24 professional actors and labeled with eight different emotions (neutral, calm, happy, sad, angry, fearful, disgust, surprised).
-  • 2.
-    CREMA-D [Cao et al., 2014]: It is an English dataset of 7442 videos recorded by 91 actors.
-Actors spoke from a selection of 12 sentences.
-The sentences used one of six emotions (anger, disgust, fear, happy, neutral, and sad) with four different intensities (low, medium, high, and unspecified).
-We selected these databases as they are commonly used for the emotion recognition task, and their raw data is accessible.
-We performed 6-fold and 10-fold cross-validation for the RAVDESS and CREAM-D datasets to ensure a fair comparison with previous works.
-The speaker identities are different in the fine-tuning and evaluation phases.
+Fig. 4: Quantitative results of the audio reconstruction (a) and visual reconstruction (b) using the VQ-MAE-AV model.
 (a) Peak Signal-to-Noise Ratio (PSNR in dB) on the y-axis as a function of masking ratio (%) on the x-axis.
 (b) Signal-to-Distortion Ratio (SDR in dB) on the y-axis as a function of masking ratio (%) on the x-axis.
-Figure 4: Quantitative results of the visual reconstruction (a) and audio reconstruction (b).
+Fig.
+5: Quantitative results of the visual reconstruction (a) and audio reconstruction (b).
 The line represents the mean value across the test dataset, while the shaded area depicts the standard deviation.
@@
 VQ-VAE architectures
-The VQ-VAE-audio (respectively VQ-VAE-visual) architecture is symmetrical concerning the encoder and the decoder, with three 1D (respectively 2D) convolution for the encoder or transposed convolution for the decoder layers and a residual convolution layer.
+The VQ-VAE-A (respectively VQ-VAE-V) architecture is symmetrical concerning the encoder and the decoder, with three 1D (respectively 2D) convolution for the encoder or transposed convolution for the decoder layers and a residual convolution layer.
 The VQ-VAE models process each frame independently with no time dependency.
-For each speech power spectrogram frame of size $D=513$, the VQ-VAE-audio encoder compresses it into a discrete latent vector (a column of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$) of size $D^{\prime}=64$.
-For each image frame of size $(H=96,W=96,C=3)$, the VQ-VAE-visual encoder compresses it into a discrete latent representation of size $(H^{\prime}=24,W^{\prime}=24)$.
-The VQ-VAE-audio and the VQ-VAE-visual codebooks contain, respectively, $k_{a}=256$ and $k_{v}=512$ codes of dimension $e_{a}=8$ and $e_{v}=4$.
-Such a low dimension is chosen to increase the use of the different codes in the codebook [Yu et al., 2021].
-Table 1: Accuracy (%) and F1 score (%) results of audiovisual SER.
-VQ-MAE-AV is pre-trained with both the generative (3.6.1) and contrastive (3.6.2) loss functions using the Query2Emo pooling strategy.
-The best scores are in bold, and the second-best scores are underlined.
-RAVDESS CREMA-D
-Method Accuracy F1 score Method Accuracy F1 score
-AV-LSTM [Ghaleb et al., 2019] 65.80 - AV-LSTM [Ghaleb et al., 2019] 72.90 -
-MuLT [Tsai et al., 2019] 76.60 77.30 MATER [Ghaleb et al., 2020] 67.20 -
-CFN-SR [Fu et al., 2021] 75.76 - RAVER [Goncalves & Busso, 2022] 77.30 -
-MATER [Ghaleb et al., 2020] 76.30 - MulT Base [Tran & Soleymani, 2022] 68.87 -
-AVT [Chumachenko et al., 2022] 79.20 78.20 MulT Large [Tran & Soleymani, 2022] 70.22 -
-MDVAE [Sadok et al., 2024] 79.30 80.70 AV-Gating [Ghaleb et al., 2019] 74.00 -
-VQ-MAE-AV (ours) 84.80 84.50 VQ-MAE-AV (ours) 80.40 80.00
-Table 2: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the attention pooling strategy for the fine-tuning.
-‘Pre-train’ refers to the training of the VQ-MAE-AV for the unmasking task on the VoxCeleb2 database.
-‘Freeze’ refers to the freezing of the VQ-MAE-AV encoder.
-Method Pre-train Freeze Accuracy (%)
-VQ-MAE-AV 29.6
-VQ-MAE-AV 70.5
-VQ-MAE-AV 81.5
-Table 3: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, fine-tuned on RAVDESS for different encoder depths.
-Method Parm. (M) Acc. (%) F1 score (%)
-VQ-MAE-AV-6 8.5 75.7 76.0
-VQ-MAE-AV-12 13.5 81.5 80.1
-VQ-MAE-AV-16 16.8 82.4 82.4
-VQ-MAE-AV-20 20.2 81.3 81.3
-Table 4: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, and using the Query2Emo pooling strategy for the fine-tuning.
-‘Generative’ corresponds to the loss function in Eq.
-3.6.1 and ‘Contrastive’ corresponds to the loss function in Eq.
-3.6.2.
-Method Contrastive Generative Accuracy (%)
-VQ-MAE-AV 75.2
-VQ-MAE-AV 84.3
-VQ-MAE-AV 84.8
-Table 5: Performance of VQ-MAE-AV without contrastive learning for different encoder and decoder architectures, using the attention pooling strategy for the fine-tuning.
-SAF stands for self-attention fusion and CAF stands for cross-attention fusion.
-Method Param. (M) Encoder Decoder Acc. (%)
-VQ-MAE-AV 13.5 SAF SAF 81.5
-VQ-MAE-AV 25.0 CAF SAF 82.8
-VQ-MAE-AV 18.4 SAF CAF 81.8
-VQ-MAE-AV 30.0 CAF CAF 83.0
-Table 6: Performance of VQ-MAE-AV using the self-attention fusion strategy for both encoder and decoder, without contrastive learning, and using different pooling strategies for the fine-tuning.
-Pooling strategy Accuracy (%) F1 score (%)
-Mean Pooling 78.1 78.4
-Attention Pooling 81.5 80.1
-Query2Emo 84.3 84.8
-Table 7: Comparison of performance between unimodal VQ-MAE-A-12 or VQ-MAE-V-12 and multimodal VQ-MAE-AV-12 (using SAF for both encoder and decoder and without contrastive learning).
-Method Accuracy (%) F1 score (%)
-VQ-MAE-A-12 73.2 72.8
-VQ-MAE-V-12 74.1 73.9
-VQ-MAE-AV-12 81.5 80.1
+For each speech power spectrogram frame of size $D=513$, the VQ-VAE-A encoder compresses it into a discrete latent vector (a column of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$) of size $D^{\prime}=64$.
+For each image frame of size $(H=96,W=96,C=3)$, the VQ-VAE-V encoder compresses it into a discrete latent representation of size $(H^{\prime}=24,W^{\prime}=24)$.
+The VQ-VAE-A and the VQ-VAE-V codebooks contain, respectively, $k_{a}=256$ and $k_{v}=512$ codes of dimension $e_{a}=8$ and $e_{v}=4$.
+Such a low dimension is chosen to increase the use of the different codes in the codebook (Yu et al., 2021).
 VQ-MAE-AV architectures
-The VQ-MAE-AV model uses $L=12$ attention blocks in the encoder and $L^{\prime}=4$ in the decoder.
-Each self-attention layer of a block is divided into four heads.
+The VQ-MAE-AV model employs an encoder comprising $L$ attention blocks, where $L$ is set to 12 unless otherwise specified, and a decoder featuring 4 attention blocks.
+Each self-attention layer of a block is divided into 4 heads.
 By default, the parameters of the discrete audio and visual tokens ($d$, $h$, and $w$) are set to 4.
@@
 Self-supervised training details
-The VQ-MAE-AV is trained using the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate, with a 100-epoch warm-up period.
+The VQ-MAE-AV is trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with a cosine scheduler to adjust the learning rate, with a 100-epoch warm-up period.
 The parameters of the optimizer, similar to He et al.
-[2022], are $\beta_{2}=0.9$, $\beta_{2}=0.95$, and weight_decay$=0.05$.
-The base learning rate follows the linear scaling rule [Goyal et al., 2017] $lr=(\texttt{base\_lr}=1e-3)\times(\texttt{batchsize}=128)/256$.
+(2022), are $\beta_{2}=0.9$, $\beta_{2}=0.95$, and weight_decay$=0.05$.
+The base learning rate follows the linear scaling rule (Goyal et al., 2017) $lr=(\texttt{base\_lr}=1e-3)\times(\texttt{batchsize}=128)/256$.
 We distributed the pre-training of VQ-MAE-AV on 4 NVIDIA HGX A100.
@@
 Fine-tuning details
-For the fine-tuning process, we also use the AdamW optimizer [Loshchilov & Hutter, 2017] with a cosine scheduler to adjust the learning rate and with a 40-epoch warm-up period.
+For the fine-tuning process, we also use the AdamW optimizer (Loshchilov and Hutter, 2017) with a cosine scheduler to adjust the learning rate and with a 40-epoch warm-up period.
 The parameters of the optimizer are the same as those used for the pre-training.
@@
 to reconstruct the complete audiovisual speech sequence from partially observed tokens.
-We will study the reconstruction performance for different masking ratios.
-For this experiment, we compare VQ-MAE-AV to VQ-MAE-A (A for audio) and VQ-MAE-V (V for visual), which are the unimodal versions of VQ-MAE-AV.
+We are interested in studying the reconstruction performance for different masking ratios.
+Qualitative results are presented in Fig.
+4(a) and Fig.
+4(b) for two masking ratios, 50% and 80%.
+Additional qualitative results are available online (see the webpage address at the end of Section 1).
+In this experiment, we compare VQ-MAE-AV to VQ-MAE-A (A for audio) and VQ-MAE-V (V for visual), which are the unimodal versions of VQ-MAE-AV.
 The average quality performance for the speech and visual modalities is evaluated using the VoxCeleb2 test set.
@@
 Fig.
-4(a) and Fig.4(b) present the PSNR and SDR curves, respectively, for the reconstruction quality of the visual and audio modalities as a function of the masking ratio.
+5(a) and Fig.5(b) present the PSNR and SDR curves, respectively, for the reconstruction quality of the visual and audio modalities as a function of the masking ratio.
 Notably, VQ-MAE-AV outperforms VQ-MAE-V for masking ratios greater than $50$%.
@@
 4.3 Audiovisual emotion recognition
-Table 1 compares the emotion recognition performance (accuracy and F1 score) of the proposed VQ-MAE-AV model (using the cross-attention fusion strategy for both the encoder and decoder) with the performance of several audiovisual state-of-the-art methods.
-The experimental comparison includes LSTM-based methods [Ghaleb et al., 2019] and Transformer-based methods [Tsai et al., 2019, Chumachenko et al., 2022, Goncalves & Busso, 2022].
-MATER [Ghaleb et al., 2020] uses distinct Transformer architectures for each modality, merging the embeddings via mean pooling.
-CFN-SR [Fu et al., 2021] introduces a cross-modality learning approach, employing self-attention and residual structures to model inter- and intra-modality interactions.
-We also consider the AVT method [Chumachenko et al., 2022], which uses self-attention fusion and modality dropout to address the challenge of one modality being absent.
-Additionally, RAVER [Goncalves & Busso, 2022] addresses challenges related to modality alignment, temporal information capture, and handling missing features.
-The experimental comparison also includes unsupervised audiovisual approaches, such as the multimodal dynamical VAE (MDVAE) [Sadok et al., 2024], which employs a hierarchical latent space to segregate static from dynamic information and modality-common from modality-specific information.
-Additionally, we compare with MulT [Tran & Soleymani, 2022], a self-supervised approach using the Transformer architecture and pre-trained on the VoxCeleb dataset with a pretext task of reconstructing masked frames (with a masking ratio of 15%).
-As can be seen in Table 1, the VQ-MAE-AV model with Query2Emo outperforms the most recent methods overall.
-VQ-MAE-AV achieves $9.04$%, $5.60$% and $5.50$% better accuracy than the CFN-SR [Fu et al., 2021], AVT [Chumachenko et al., 2022] and MDVAE [Sadok et al., 2024] methods for the RAVDESS dataset, respectively.
-Regarding the CREMA-D dataset, VQ-MAE-AV achieves $7.50$%, $6.40$%, and $3.10$% better accuracy than the AV-LSTM [Ghaleb et al., 2019], AV-Gating [Ghaleb et al., 2019], and RAVER [Goncalves & Busso, 2022].
+Table 1: Audiovisual emotion recognition results in terms of accuracy (%) and F1 score (%) on the RAVDESS and CREMA-D datasets
+.
+The best scores are in bold, and the second-best scores are underlined.
+RAVDESS CREMA-D Method Accuracy F1 score Method Accuracy F1 score AV-LSTM (Ghaleb et al., 2019) 65.80 - AV-LSTM (Ghaleb et al., 2019) 72.90 - MuLT (Tsai et al., 2019) 76.60 77.30 MATER (Ghaleb et al., 2020) 67.20 - CFN-SR (Fu et al., 2021) 75.76 - AV-Gating (Ghaleb et al., 2019) 74.00 - MATER (Ghaleb et al., 2020) 76.30 - MulT Base (Tran and Soleymani, 2022) 68.87 - AVT (Chumachenko et al., 2022) 79.20 78.20 MulT Large (Tran and Soleymani, 2022) 70.22 - MDVAE (Sadok et al., 2024) 79.30 80.70 RAVER (Goncalves and Busso, 2022) 77.30 - VQ-MAE-AV (ours) 84.80 84.50 VQ-MAE-AV (ours) 80.40 80.00
+Table 2: Accuracy (%) and F1 score (%) results on DFEW and Aff-Wild2. The best scores are in bold, and the second-best scores are underlined.
+DFEW Aff-Wild2
+Method Accuracy F1 score Method Accuracy F1 score
+C3D+LSTM (Zhang et al., 2023) 65.17 - CT-VIPL (Liu et al., 2020) 64.00 33.00
+ResNet-18+LSTM (Zhang et al., 2023) 64.32 - Multi-modal ER (Jin et al., 2021) 63.00 40.20
+T-MEP (Zhang et al., 2023) 68.85 - CNN-RNN (Antoniadis et al., 2021) 66.80 55.50
+VQ-MAE-AV (ours) 70.30 69.80 VQ-MAE-AV (ours) 69.70 68.90
+In this section, we evaluate the emotion recognition performance of the proposed VQ-MAE-AV model on 4 different datasets, including 2 recorded in the wild, and we compare the model’s performance with that of 15 state-of-the-art methods.
+The performance is measured in terms of accuracy and F1 score.
+For this experimental comparison, we chose the best configuration of the VQ-MAE-AV model, which uses the cross-attention fusion strategy for the encoder and decoder, both generative and contrastive loss functions for self-supervised pre-training on VoxCeleb2, and the Query2Emo strategy for supervised fine-tuning.
+Other configurations will be investigated in the ablation study of Section 4.4.
+4.3.1 Results in controlled conditions
+We start by discussing the emotion recognition results on RAVDESS and CREMA-D, the two datasets recorded in controlled lab conditions.
+The experimental comparison includes LSTM-based methods (Ghaleb et al., 2019) and Transformer-based methods (Tsai et al., 2019; Chumachenko et al., 2022; Goncalves and Busso, 2022).
+MATER (Ghaleb et al., 2020) uses distinct Transformer architectures for each modality, merging the embeddings via mean pooling.
+CFN-SR (Fu et al., 2021) introduces a cross-modality learning approach, employing self-attention and residual structures to model inter- and intra-modality interactions.
+We also consider the AVT method (Chumachenko et al., 2022), which uses self-attention fusion and modality dropout to address the challenge of one modality being absent.
+Additionally, RAVER (Goncalves and Busso, 2022) addresses challenges related to modality alignment, temporal information capture, and handling missing features.
+The experimental comparison also includes the multimodal dynamical VAE (MDVAE) (Sadok et al., 2024), which employs an unsupervised hierarchical latent representation to segregate static from dynamic information and modality-common from modality-specific information.
+Additionally, we compare with MulT (Tran and Soleymani, 2022), a self-supervised approach using the Transformer architecture and pre-trained on the VoxCeleb dataset with a pretext task of reconstructing masked frames (with a masking ratio of 15%).
+As can be seen in Table 1, the VQ-MAE-AV model outperforms the most recent methods overall.
+VQ-MAE-AV achieves $9.04$%, $5.60$% and $5.50$% better accuracy than the CFN-SR (Fu et al., 2021), AVT (Chumachenko et al., 2022) and MDVAE (Sadok et al., 2024) methods for the RAVDESS dataset, respectively.
+Regarding the CREMA-D dataset, VQ-MAE-AV achieves $7.50$%, $6.40$%, and $3.10$% better accuracy than the AV-LSTM (Ghaleb et al., 2019), AV-Gating (Ghaleb et al., 2019), and RAVER (Goncalves and Busso, 2022).
+4.3.2 Results on in-the-wild datasets
+We now discuss the emotion recognition results of the proposed approach in more realistic recording conditions, using the two in-the-wild datasets DFEW and Aff-Wild2.
+The experimental comparison includes methods based on combinations of convolutional and LSTM networks (Liu et al., 2020; Jin et al., 2021; Zhang et al., 2023; Antoniadis et al., 2021) and a Transformer-based method (Zhang et al., 2023).
+As can be seen in Table 2, the VQ-MAE-AV model with Query2Emo achieves superior performance compared to recent methods on the DFEW dataset.
+Specifically, VQ-MAE-AV outperforms the Transformer-based method T-MEP (Zhang et al., 2023) by $1.45\%$ in accuracy.
+On the Aff-Wild2 dataset, VQ-MAE-AV achieves a $2.9\%$ improvement in accuracy and a $13.4\%$ boost in F1 score compared to the CNN-RNN model of Antoniadis et al.
+(2021).
+These results confirm that the VQ-MAE-AV model also performs well on real audiovisual speech data recorded in uncontrolled environments.
 Overall, the experimental results demonstrate the effectiveness of the proposed audiovisual self-supervised representation learning technique for SER.
-This indicates that VQ-MAE-AV learns audiovisual representations that are effectively transferable to the emotion recognition task, resulting in improved performance compared to audiovisual state-of-the-art methods.
-4.4 Ablation study and model properties
-We conduct a series of experiments to evaluate the impact of various hyperparameters and model designs on the emotion recognition performance of the proposed VQ-MAE-AV model.
-In the following paragraphs, we will present and discuss the findings of these experiments.
-All the ablation study is done on RAVDESS.
+This indicates that VQ-MAE-AV learns audiovisual representations that are effectively transferable to the emotion recognition task, resulting in improved performance compared to state-of-the-art methods.
+Table 3: Performance of the baseline VQ-MAE-AV model (first row, in blue), compared to its performance when the encoder is frozen during supervised training on RAVDESS (second row) or when self-supervised pre-training on VoxCeleb2 is discarded (third row).
+Pre-training Frozen encoder Accuracy (%)
+81.5
+70.5
+29.6
+Table 4: Performance, number of parameters (in millions), and floating-point operations (FLOPs, in billions) of the baseline VQ-MAE-AV model (second row, in blue), compared to variations where the number $L$ of attention blocks in the encoder is modified (other rows).
+Encoder depth Param. (M) FLOPs (G) runtime (ms) Acc. (%) F1 score (%)
+$L=6$ 8.5 5.9 0.37 75.7 76.0
+$L=12$ 13.5 9.4 0.67 81.5 80.1
+$L=16$ 16.8 11.7 0.86 82.4 82.4
+$L=20$ 20.2 14.1 0.98 81.3 81.3
+Table 5: Performance of the VQ-MAE-AV model when using the contrastive (1st row), the generative (2nd row), or both (3rd row) loss functions during the self-supervised pre-training.
+Contrastive Generative Accuracy (%)
+75.2
+84.3
+84.8
+Table 6: Performance and number of parameters of the baseline VQ-MAE-AV model (first row, in blue), compared to variations with different modality-fusion strategies at the encoder and decoder (other rows).
+Encoder Decoder Param. (M) Acc. (%)
+SAF SAF 13.5 81.5
+CAF SAF 25.0 82.8
+SAF CAF 18.4 81.8
+CAF CAF 30.0 83.0
+Table 7: Performance of the baseline VQ-MAE-AV model (1st row, in blue), compared to variations of the emotion recognition model (other rows).
+Emotion recognition model Accuracy (%) F1 score (%)
+Attention Pooling 81.5 80.1
+Mean Pooling 78.1 78.4
+Query2Emo 84.3 84.8
+Table 8: Performance of the baseline VQ-MAE-AV model using the audio and visual modalities (1st row, in blue), compared to the equivalent model using only the audio modality (2nd row) or the visual modality (3rd row).
+Modality Accuracy (%) F1 score (%)
+Audio + visual 81.5 80.1
+Audio 73.2 72.8
+Visual 74.1 73.9
+4.4 Ablation study
+In this section, we present a series of ablation experiments using the RAVDESS dataset.
+Our objective is to evaluate the impact of various hyperparameters and model designs on the emotion recognition performance of the proposed VQ-MAE-AV model.
+In the ablation experiments, we systematically vary one single hyperparameter or model block at a time, starting from a baseline configuration.
+This baseline uses self-attention fusion for both the encoder and the decoder (see Section 3.5), it uses only the generative cross-entropy loss function for self-supervised pre-training (see Section 3.6), and it relies on attention pooling for supervised fine-tuning (see Section 3.7).
+This configuration differs from the best-performing one used for the audiovisual emotion recognition experiments presented in Section 4.3, which used the cross-attention fusion strategy for the encoder and decoder, both generative and contrastive loss functions for self-supervised pre-training, and the Query2Emo strategy for supervised fine-tuning.
+For clarity and ease of interpretation, the baseline configuration is highlighted in blue in the tables of results of the present section.
 4.4.1 Impact of pre-training and fine-tuning
 Table 8 shows the significance of pre-training and fine-tuning the VQ-MAE-AV model for audiovisual SER.
-Pre-training the model for the unmasking task on the VoxCeleb2 database substantially improves the emotion recognition performance, with the accuracy rising from $29.6$% to $81.5$%.
-Fine-tuning the encoder is also essential, as keeping it frozen leads to a $11$% drop in accuracy.
+Self-supervised pre-training of the model for the unmasking task on the VoxCeleb2 dataset substantially improves the emotion recognition performance, with the accuracy rising from $29.6$% to $81.5$%.
+Fine-tuning the encoder during the supervised training on RAVDESS is also essential, as keeping it frozen leads to a $11$% drop in accuracy.
 4.4.2 Impact of the encoder depth
-Table 7 shows the impact of encoder depth on emotion recognition performance by varying the number of attention blocks ($L$), where a model with n attention blocks in the encoder will be denoted by VQ-MAE-AV-n.
-The results show that increasing the number of blocks in the encoder leads to improved performance to a certain extent.
-However, when the number of blocks becomes too high (as in the case of VQ-MAE-AV-20), there is no further improvement, and performance actually decreases by $1.1$% accuracy compared to VQ-MAE-AV-16.
-4.4.3 Impact of the contrastive learning
-The influence of contrastive learning on emotion recognition is illustrated in Table 7.
-When training VQ-MAE-AV exclusively with the contrastive loss, the achieved accuracy is $75.2$%, which is $9.1$% lower than when using only the generative loss.
-Notably, the combination of both losses, contrastive and generative, leads to improved emotion recognition, resulting in a $0.5$% accuracy boost compared to using the generative loss alone.
-4.4.4 Impact of the encoder/decoder architecture
-Table 7 shows the performance of the four different configurations of the VQ-MAE-AV model described in Section 3.5, including all possible combinations of the fusion strategies for the VQ-MAE-AV encoder and decoder: self-attention fusion for both the encoder and decoder (SAF-SAF); self-attention fusion for the encoder and cross-attention fusion for the decoder (SAF-CAF); cross-attention fusion for the encoder and self-attention fusion for the decoder (CAF-SAF); and cross-attention fusion for both the encoder and decoder (CAF-CAF).
-Among these configurations, CAF-CAF achieves the highest accuracy with a $1.5$% improvement over SAF-SAF, followed by CAF-SAF with a $1.3$% improvement over SAF-SAF, and then SAF-CAF with only a $0.3$% improvement.
-The cross-attention fusion encoder architecture achieves the best performance in emotion recognition, as shown by the CAF-CAF and CAF-SAF configurations.
+Table 8 shows the impact of varying the number $L$ of attention blocks in the VQ-MAE-AV encoder, in terms of emotion recognition performance (accuracy and F1 score), number of parameters, the number of floating-point operations (FLOPs) and the runtime (ms) required to make a forward pass in the model for 2 second-long input sequence.
+The results indicate that increasing the number of blocks in the encoder leads to improved emotion recognition performance up to $L=16$; the accuracy decreases by $1.1$% for $L=20$.
+Moreover, it can be seen that this increase in performance comes at the expense of a larger model that requires more FLOPs to compute the prediction.
+4.4.3 Impact of the generative and contrastive loss functions
+Table 8 shows the impact of using either the generative, the contrastive, or both loss functions during the self-supervised pre-training of the model.
+This experiment is conducted from the best-performing VQ-MAE-AV model (corresponding to the results in Section 4.3) instead of the baseline one.
+It can be seen that the best performance is achieved when using the two loss functions.
+When training VQ-MAE-AV exclusively with the contrastive (resp.
+generative) loss, the accuracy drops by $9.6$% (resp.
+$0.5$%).
+4.4.4 Impact of the modality-fusion strategies at the encoder and decoder
+Table 8 shows the the performance obtained with the different modality-fusion strategies described in Section 3.5, including all possible combinations for the VQ-MAE-AV encoder and decoder: self-attention fusion for both the encoder and decoder (SAF-SAF); self-attention fusion for the encoder and cross-attention fusion for the decoder (SAF-CAF); cross-attention fusion for the encoder and self-attention fusion for the decoder (CAF-SAF); and cross-attention fusion for both the encoder and decoder (CAF-CAF).
+As can be seen, CAF-CAF achieves the highest accuracy with a $1.5$% improvement over SAF-SAF, followed by CAF-SAF with a $1.3$% improvement over SAF-SAF, and then SAF-CAF with only a $0.3$% improvement.
+The cross-attention fusion architecture at the encoder achieves the best performance, as shown by the CAF-CAF and CAF-SAF configurations.
 However, there exists a trade-off between performance and the number of model parameters.
 Notably, CAF-CAF involves slightly more than twice as many parameters as SAF-SAF.
-4.4.5 Impact of the pooling strategy
-Table 7 presents the impact of various pooling strategies when fine-tuning the model for audiovisual SER.
-The results reveal that attention-based pooling techniques, such as attention pooling and Query2Emo, outperform mean pooling.
-Among the two attention-based methods, Query2Emo outperforms attention pooling, with an accuracy gain of $2.8$%.
+4.4.5 Impact of the emotion recognition model
+Table 8 illustrates the impact of various emotion recognition models used on top of the VQ-MAE-AV encoder for supervised SER.
+As could be expected, attention pooling and Query2Emo outperform the naive mean pooling strategy.
+Among the two, Query2Emo performs the best, with an accuracy gain of $2.8$%.
+4.4.6 Impact of the modalities
+Table 8 compares the performance of the proposed multimodal model with the equivalent model trained using only the audio or visual modalities.
+It can be seen that exploiting both modalities greatly improves the performance, with accuracy gains of $5.6$% and $8.3$% compared to using only the visual and audio modalities, respectively.
 4.4.7 Impact of the audio and visual discrete token size
 Fig.
-5 shows the impact of the dimensions of the discrete tokens for the visual ($h$ and $w$) and audio ($d$) modalities on the audiovisual SER performance.
-The values of $h$ and $w$ represent the visual token size on the horizontal and vertical axes, respectively, while $d$ represents the token size on the frequency axis of the audio modality.
-Our study reveals that the performance of emotion recognition is impacted by both ($h,w$) and $d$ values.
-Therefore, it is important to select these parameters carefully.
+6 shows the impact of the dimension of the discrete visual and audio tokens in terms of emotion recognition accuracy.
+In this figure, $h$ and $w$ represent the horizontal and vertical dimensions of the token in the visual modality, while $d$ represents the frequency dimension of the tokens in the audio modality.
+Moreover, we only consider the case $h=w$.
+The results reveal that it is important to select these parameters carefully, as their value significantly impacts the performance.
 Based on our experiments, we recommend fixing them to ($h=w=4,~{}d=4$).
-Figure 5: Impact of the discrete audio and visual token size on emotion recognition.
-4.4.7 Impact of the audiovisual inputs
-Table 7 compares our approach VQ-MAE-AV with its unimodal versions (VQ-MAE-V and VQ-MAE-A).
-The VQ-MAE-AV model exhibits improved performance over VQ-MAE-V and VQ-MAE-A, with gains of $5.6$% and $8.3$% in accuracy, respectively.
-This indicates the effectiveness of incorporating audiovisual information in the self-supervised paradigm, leading to better performance in emotion recognition.
+Fig. 6: Impact of the discrete audio and visual token size on emotion recognition.
+4.4.8 Summary
+The ablation study highlighted key factors that impact SER performance, including attention blocks, model structure, hyperparameters, and training losses.
+Fine-tuning on emotional datasets proved crucial, as VQ-MAE-AV initially learns general representations through masked token reconstruction, requiring adaptation for SER (see Table 8).
+Integrating both audio and visual modalities further enhanced performance over single-modality inputs (see Table 8).
+Additionally, increasing attention blocks in the encoder improves SER performance up to a saturation point (see Table 8).
+Other factors, such as cross-attention fusion, token size, contrastive learning, and emotion recognition models, offer smaller but notable improvements.
 5 Conclusion
-Masked autoencoder modeling is a versatile self-supervised learning approach that can be adapted to various types of data.
+Masked autoencoding is a versatile self-supervised learning approach that can be adapted to various types of data.
 This paper introduced the VQ-MAE-AV model for learning representations of audiovisual speech data, which could be extended to other multimodal sequential data.
@@
 These representations were then divided into multiple discrete tokens, with spatio-temporal tokens for the visual modality and spectro-temporal tokens for the audio modality.
-Pre-trained on the VoxCeleb2 dataset and fine-tuned on standard emotional audiovisual speech datasets, the experiments showed that the VQ-MAE-AV model effectively combined the audio and visual modalities for audiovisual SER, outperforming several state-of-the-art methods.
-For future work, we plan to investigate other masking strategies and increase the resolution of the images to further improve emotion recognition performance.
-We could also explore using the representation learned by the proposed VQ-MAE-AV model for other downstream tasks in audiovisual speech processing, such as speaker recognition.
+Pre-trained on the VoxCeleb2 dataset and fine-tuned on emotional audiovisual speech datasets, the experiments showed that the VQ-MAE-AV model effectively combines the audio and visual modalities for audiovisual SER, outperforming several state-of-the-art methods in both controlled and in-the-wild conditions.
+During self-supervised pre-training, the VQ-MAE-AV model learns general-purpose audiovisual speech representations that require fine-tuning for SER.
+Unlike contrastive SSL approaches, where simple linear probing often yields satisfactory results, our model necessitates fine-tuning on emotional datasets to effectively adapt its representations for emotion recognition.
+This is a limitation, but it is also an opportunity, as the learned representations could be adapted for various tasks beyond SER (Baevski et al., 2022), such as audiovisual speech recognition, speaker identification, or cross-modal generative tasks, like in AnCoGen (Sadok et al., 2025).
+VQ-MAE-AV employs masked modeling as an SSL pre-training paradigm, enabling efficient cross-modal integration.
+This approach allows the model to learn general audiovisual representations without relying on complex fusion mechanisms, making it adaptable to incorporate additional modalities such as human gesture and pose.
+This is a particularly interesting avenue for future work on multimodal human behavior analysis, in particular emotion recognition.
 Acknowledgments
-This work was performed using computational resources from the “Mésocentre” computing center of Université Paris-Saclay, CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-saclay.fr/)
+This work was supported by Randstad corporate research chair.
+It was performed using computational resources from the “Mésocentre” computing center of Université Paris-Saclay, CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-saclay.fr/).